### v1

# GSPRec: Temporal-Aware Graph Spectral Filtering for Recommendation

Abstract
Graph-based recommendation systems are effective at modeling collaborative patterns but often suffer from two limitations: overreliance on low-pass filtering, which suppresses user-specific signals, and omission of sequential dynamics in graph construction. We introduce GSPRec, a graph spectral model that integrates temporal transitions through sequentially-informed graph construction and applies frequency-aware filtering in the spectral domain. GSPRec encodes item transitions via multi-hop diffusion to enable the use of symmetric Laplacians for spectral processing. To capture user preferences, we design a dual-filtering mechanism: a Gaussian bandpass filter to extract mid-frequency, user-level patterns, and a low-pass filter to retain global trends. Extensive experiments on four public datasets show that GSPRec consistently outperforms baselines, with an average improvement of 6.77% in NDCG@10. Ablation studies show the complementary benefits of both sequential graph augmentation and bandpass filtering.
1 Introduction
Graph-based recommender systems are constrained by two core problems: using low-pass spectral filtering often suppresses distinct user signals in favor of broad popularity patterns, and the order of user interactions is typically neglected in their graph structures. Therefore, recommender systems should leverage both collaborative patterns and the rich temporal signals embedded within these interaction sequences [19]. Graph-based Collaborative Filtering (CF) methods [14, 42, 49] have achieved strong performance, but they face two limitations. First, models such as LightGCN [14], which rely on linear embedding propagation, act as low-pass filters [42, 33] by smoothing signals across the user-item graph, which encourages popularity but obscures user-specific patterns critical for personalization [3]. Second, while some GCN-based recommenders incorporate temporal data, they do so via separate sequential modeling components [26, 16] or by methods that do not integrate order into the graph structure for spectral analysis, which means they still discard sequential information needed to distinguish user preferences from popularity patterns [25, 42, 52, 38, 55].
The filtering behavior inherent in many GCNs motivates a closer look through the lens of Graph Signal Processing (GSP), which generalizes classical signal processing to data defined on irregular graph structures [43, 34]. A core technique in GSP is spectral filtering, which transforms graph signals using the eigenbasis of the graph Laplacian [20]. Lowpass filters, which retain only low-frequency components, are widely adopted under the assumption that graph signals are typically smooth [43, 42, 14]. However, mid- and high-frequency components often encode task-relevant, localized, or personalized variations [7, 46, 17] motivating the design of spectral filters that can selectively amplify non-smooth components. Despite GSP’s potential, its adoption in recommender systems [25, 53, 52, 42] has largely focused on lowpass filtering to capture popular preference patterns, overlooking user-specific patterns.
To overcome these limitations, we introduce GSPRec, a novel GSP-based framework. GSPRec achieves temporal awareness in spectral recommenders through two components: First, its sequentially-aware graph construction encodes multi-hop item transition patterns from user histories via diffusion into a symmetric graph structure, making temporal dynamics an intrinsic property for spectral analysis. Second, this graph empowers a frequency-aware dual-filtering, featuring a Gaussian bandpass filter to isolate vital mid-frequency user-specific preferences, complemented by a low-pass filter for popularity patterns.
In this work, we encode the interaction graph with item-to-item transitions derived from user interaction sequences. A multi-hop diffusion process with exponential decay captures sequential dependencies while preserving the benefits of symmetric graph structures. These symmetric structures ensure real eigenvalues and orthogonal eigenvectors [6], which are crucial for stable spectral decomposition and well-defined filtering operations [34]. Unlike prior work that infers item-item similarity from collaborative information [53], our approach aggregates explicit sequential signals into the graph topology. Figure 1 illustrates this shift from sequence-based to graph-based modeling.
Moreover, we design a Gaussian bandpass filter tailored to recommendation graphs. Informed by empirical spectral energy distributions [38, 46], the filter selectively captures mid-frequency components that are often associated with user-specific and sequential patterns positioned between low-frequency popularity signals and high-frequency noise. The Gaussian profile provides optimal localization in both spectral and vertex domains [43, 34], which enables smooth transitions across frequencies and reduces ringing artifacts commonly introduced by sharp filter cutoffs [12]. We summarize our contributions as follows:
(a) Sequential pattern
(b) Bidirectional pattern
(c) Undirected graph
Figure 1: (a) Unidirectional transitions in models like RNNs [15, 18] process items in strict temporal order. (b) Bidirectional models such as BERT4Rec [45] allow context flow in both directions, but only between adjacent items. (c) Undirected item graph from user sequences. Solid lines show the initial symmetric graph $\mathbf{S}^{\prime}$. Orange dashed lines indicate higher-order relationships learned via multi-hop diffusion (Eq. 2) on $\mathbf{S}^{\prime}$, which aggregates weighted sequential paths ($\alpha$ and $d$). The strength of sequence-based connections between items, derived from the aggregated weighted paths of the diffusion process, is reflected in these symmetric links, making the graph ready for spectral filtering. This sequence is a running example to illustrate graph construction and transition modeling.
  • We propose a novel graph construction method that first derives a symmetric item connectivity graph from user sequences, then encodes higher-order sequential dependencies via multi-hop diffusion on this symmetric structure.
  • We propose a Gaussian bandpass filter that targets mid-frequency components.
  • We provide theoretical analysis establishing the stability and convergence of our sequential diffusion and filtering mechanisms.
  • We validate our framework on four public benchmarks, showing that integrating sequential transitions with mid-frequency filtering leads to consistent performance gains over baselines.
GSPRec achieves average gains of 6.77% in NDCG@10 and 14.88% in MRR@10 across four public benchmark datasets, consistently outperforming state-of-the-art baselines. These results highlight the effectiveness of our two core design choices: sequential graph encoding and spectral filtering.
2 Related Work & Background
Table 1: Comparison with existing popular methods.
Method Task Spectral Filtering Temporal Info. Graph Structure Message Passing Mid-Freq
Message Passing GCNs:
NGCF [49] General User-Item
LightGCN [14] General User-Item
LR-GCCF [4] General User-Item
IMP-GCN [23] General User-Item
Spectral Filtering Methods (GSP-based):
GF-CF [42] General LP User-Item
PGSP [25] General LP User-Item
HiGSP [53] General LP User-Item
FaGSP [52] General LP User-Item
Non-Propagative / Simplified Methods:
SimpleX [28] General User-Item
UltraGCN [29] General ApproxLP User-Item
Sequential Recommendation:
BERT4Rec [45] Sequential Sequence
SASRec [18] Sequential Sequence
GSPRec (Ours) General BP+LP UI+Seq
Graph-Based Collaborative Filtering:
Graph-based collaborative filtering represents users and items as nodes in a bipartite graph, modeling interactions through various approaches [14, 49, 55, 60, 17, 59]. Recent works address challenges including over-smoothing [3, 42], popularity bias [58], representation collapse [36], fairness [2, 56], and efficiency [22, 51], while others improve representations through disentanglement [50, 39, 9] or architectural simplification [29, 28]. Temporal dynamics have been incorporated through sequential models [18, 45, 21] or temporal graphs [57, 35]. Our approach encodes temporal locality directly as edge-level structure and applies spectral filtering without message passing.
Background on Graph Signal Processing:
Graph Signal Processing (GSP) analyzes signals on graph structures [34, 43, 40]. A graph signal maps values to nodes $\mathcal{G}=(\mathcal{V},\mathcal{E})$ as $\mathbf{x}\in\mathbb{R}^{|\mathcal{V}|}$. The normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ decomposes into eigenvalues $\boldsymbol{\Lambda}$ and eigenvectors $\mathbf{U}$ [6], forming the graph Fourier basis. Spectral filtering transforms signals:
$\mathbf{x}^{\prime}=\mathbf{U}\,g(\boldsymbol{\Lambda})\,\mathbf{U}^{T}\mathbf{x},\quad\text{where }g(\boldsymbol{\Lambda})=\mathrm{diag}(g(\lambda_{1}),\ldots,g(\lambda_{n}))$ (1)
Low-pass filters preserve global patterns, while band-pass filters capture mid-frequency variations related to user-specific behavior [34, 12]. In our context, user preference signals and item transition patterns enable spectral analysis as an alternative to neighborhood aggregation.
Graph Signal Processing in Recommendation:
Prior work has adapted GSP to recommender systems by treating interactions as graph signals [17, 27, 25, 52, 53, 61, 38]. These approaches employ low-pass filtering for global trends, neglecting mid-frequency components where user-specific patterns often reside. Various methods integrate spectral principles with neural architectures [52, 48, 54] or explore parameter-free formulations [24]. Our approach combines band-pass and low-pass filtering with temporal structure encoding, drawing from classical GSP filter design [20, 46, 30].
3 GSPRec
Figure 2: Overview of GSPRec. Starting from user-item interaction sequences (left), we construct a unified graph joining user-item links and item-item transitions via multi-hop diffusion with exponential decay (Eq. 2). This graph is processed using two complementary spectral filters: a band-pass filter to extract mid-frequency, user-specific patterns, and a low-pass filter to capture global trends.
GSPRec consists of a sequential graph encoding procedure and a frequency-aware spectral filtering framework. These components jointly capture two complementary signal patterns in the spectral domain: (1) globally consistent preference trends across the entire user-item graph, and (2) user-specific variation patterns that distinguish individual preferences. Figure 2 summarizes the pipeline from user-item interactions to final recommendations via graph construction and spectral filtering.
3.1 Problem Formulation
Let $\mathcal{U}=\{u_{1},\ldots,u_{m}\}$ denote users and =$\mathcal{I}=\{i_{1},\ldots,i_{n}\}$ denote items. Historical interactions are triplets $\mathcal{D}=\{(u,i,t)\mid u\in\mathcal{U},i\in\mathcal{I},t\in\mathbb{R}^{+}\}$, yielding binary matrix $\mathbf{X}\in\{0,1\}^{m\times n}$ where $x_{ui}=1$ if user $u$ interacted with item $i$. For each user $u$, we define a time-ordered sequence $\mathcal{S}_{u}=[i_{1},i_{2},\ldots]$ where $(u,i_{j},t_{j})\in\mathcal{D}$ and $t_{1}<t_{2}<\cdots$. Our task is to predict future user-item interactions by jointly modeling collaborative patterns and sequential dynamics. Specifically, we aim to predict unobserved entries in $\mathbf{X}$ using a unified spectral graph approach that leverages both user preference similarities and temporal item transitions.
3.2 Recommendation Graph Construction
Our goal is to construct a unified graph that integrates collaborative filtering signals with sequential transition patterns. This allows us to employ spectral methods to model both long-term user preferences and short-term sequential behavior, which typically require symmetric graph Laplacians [6]. The core challenge is to meaningfully translate inherently directed user sequences into a symmetric item-item graph component, denoted $\tilde{\mathbf{S}}$, suitable for this unified spectral analysis.
To achieve this, we shift from traditional sequence modeling paradigms to a graph-based representation that makes temporal dynamics amenable to spectral analysis. Figure 1 illustrates this conceptual shift: unlike models that process sequences directly through (a) unidirectional or (b) locally bidirectional mechanisms, our approach transforms user interaction sequences into (c) an undirected item graph. This graph is constructed to embed sequence-derived item proximity into a symmetric structure, a critical step for enabling spectral methods that typically require such symmetry. Although sequential models excel at capturing ordered context, our graph-based formulation is designed for compatibility with spectral analysis on symmetric matrices. This enables joint analysis of collaborative and sequential signals and decomposition with well-defined spectral properties [20]. However, using directed graphs directly with spectral methods would necessitate asymmetric Laplacians [5], which have different theoretical/computational aspects beyond our current scope. To create the symmetric item-item graph component $\tilde{\mathbf{S}}$ from sequences, we follow the following process.
Initial Symmetric Sequence Graph ($\mathbf{S}^{\prime}$):
We define an initial directed transition matrix $\mathbf{S}\in\{0,1\}^{n\times n}$, where $s_{ij}=1$ if item $i$ directly precedes item $j$ in any user sequence. To adapt this for symmetric spectral methods, we form its undirected version, $\mathbf{S}^{\prime}\in\{0,1\}^{n\times n}$. In $\mathbf{S}^{\prime}$, an undirected edge $(i,j)$ exists if a transition $i\rightarrow j$ or $j\rightarrow i$ is in $\mathbf{S}$. This symmetrization is a standard practice when preparing directed graph for algorithms requiring undirected structures [47, 41]. While simplifying transition directionality, this approach allows capturing broader item relatedness from sequences, as rigid order is not always optimal and wider context can be beneficial [45]. Multi-hop diffusion (Eq. 2) operates on $\mathbf{S}^{\prime}$ to encode sequence-derived item associations.
Multi-Hop Diffusion ($\mathbf{S}^{(d)}$):
While $\mathbf{S}^{\prime}$ captures direct sequential links to model broader item influence propagated through multi-step paths within this symmetrized sequence-derived structure, we apply multi-hop diffusion directly to the symmetric graph $\mathbf{S}^{\prime}$:
$\mathbf{S}^{(d)}=\sum_{k=1}^{d}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$ (2)
where $\alpha\in(0,1)$ is an exponential decay factor that gives higher weight to shorter paths in $\mathbf{S}^{\prime}$, thus emphasizing local sequential relationships. $d$ is the diffusion depth to control the extent of influence propagation. Since $\mathbf{S}^{\prime}$ is symmetric, each power $(\mathbf{S}^{\prime})^{k}$ is symmetric, and therefore the diffused matrix $\mathbf{S}^{(d)}$ is also symmetric. The entries $S^{(d)}_{ij}$ represent a score of multi-hop, locality-aware proximity between items $i$ and $j$, learned from the aggregated sequential patterns as illustrated in Figure 3. This diffusion process is analogous to using graph diffusion kernels to define similarity [10] and converges when $\alpha<1/\rho(\mathbf{S}^{\prime})$.
Lemma 1 (Diffusion Convergence).
The diffusion process defined in Eq. 2 converges as $d\to\infty$ if $\alpha<1/\rho(\mathbf{S}^{\prime})$, where $\rho(\cdot)$ denotes the spectral radius of $\mathbf{S}^{\prime}$.
Proof.
The diffusion process as $d\to\infty$ is $\mathbf{S}^{(\infty)}=\sum_{k=1}^{\infty}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$. We can rewrite this as $\mathbf{S}^{(\infty)}=\mathbf{S}^{\prime}\sum_{j=0}^{\infty}(\alpha\mathbf{S}^{\prime})^{j}$ by letting $j=k-1$. This series converges if the standard matrix geometric series $\sum_{j=0}^{\infty}M^{j}$, where $M=\alpha\mathbf{S}^{\prime}$, converges. This occurs if the spectral radius $\rho(M)<1$, i.e., $\rho(\alpha\mathbf{S}^{\prime})<1$. Since $\rho(\alpha\mathbf{S}^{\prime})=\alpha\rho(\mathbf{S}^{\prime})$ (for $\alpha>0$), convergence requires $\alpha<1/\rho(\mathbf{S}^{\prime})$. ∎
Figure 3: Multi-hop diffusion (Eq. 2) on the symmetric base graph $\mathbf{S}^{\prime}$ (derived from initial directed sequences) with $\alpha=0.5$. Top: $\mathbf{S}^{\prime}$. Middle: Powers $(\mathbf{S}^{\prime})^{k}$. Bottom: $\mathbf{S}^{(3)}$ encoding decayed proximity.
In practice, we estimate $\rho(\mathbf{S}^{\prime})$ (e.g., via power iteration) and choose a suitable $\alpha$.
Symmetric Normalization ($\tilde{\mathbf{S}}$):
The resulting symmetric, diffused matrix $\mathbf{S}^{(d)}$ is then symmetrically normalized to produce the final item-item graph component:
$\tilde{\mathbf{S}}=\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2},\quad\text{where }\mathbf{D}_{S}=\text{diag}(\mathbf{S}^{(d)}\mathbf{1}_{n})$ (3)
This standard normalization technique [6] yields $\tilde{\mathbf{S}}$, ensuring its weights are scaled appropriately for stable spectral analysis. The weights in $\tilde{\mathbf{S}}$ reflect a transformation of user sequences into measures of item-item sequential relevance.
Interaction Normalization:
The raw user-item interaction matrix $\mathbf{X}$ is normalized to account for varying user activity levels and popularities of the items. Standard normalizations include:
$\tilde{\mathbf{X}}_{U}=\mathbf{D}_{U}^{-1/2}\mathbf{X},\quad\tilde{\mathbf{X}}_{I}=\mathbf{X}\mathbf{D}_{I}^{-1/2}$ (4)
where $\mathbf{D}_{U}=\text{diag}(\mathbf{X}\mathbf{1}_{n})$ and $\mathbf{D}_{I}=\text{diag}(\mathbf{X}^{T}\mathbf{1}_{m})$. For constructing the unified adjacency matrix $\mathbf{A}$, we use the original binary interaction matrix $\mathbf{X}$, as its structure is fundamental. The normalized versions can be used in other components, like the filter design.
Unified Graph Adjacency and Laplacian ($\mathbf{A},\mathbf{L}$):
We integrate the user-item interactions $\mathbf{X}$ and the sequence-derived item-item graph $\tilde{\mathbf{S}}$ into a single, symmetric adjacency matrix for the unified graph:
$\mathbf{A}=\begin{bmatrix}\mathbf{0}_{m\times m}&\mathbf{X}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}\end{bmatrix}$ (5)
This matrix $\mathbf{A}$ forms a heterogeneous graph where the $\mathbf{0}_{m\times m}$ block denotes no direct user-user links are explicitly modeled here, $\mathbf{X}$ captures user-item interactions, and $\tilde{\mathbf{S}}$ encodes the sequence-derived item-item relationships. From $\mathbf{A}$, we compute the symmetric normalized Laplacian, which is the cornerstone of our spectral filtering:
$\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2},\quad\text{where }\mathbf{D}=\text{diag}(\mathbf{A}\mathbf{1})$ (6)
The properties of this Laplacian are discussed in Proposition 2. Our approach, by constructing $\tilde{\mathbf{S}}$ from symmetrized sequential data and then applying diffusion, embeds sequence-derived proximity into a standard spectral framework, offering a distinct way to leverage temporal signals compared to pure sequential models that maintain strict causality [15, 18, 45].
Proposition 2 (Laplacian Validity).
The Laplacian $\mathbf{L}$, constructed from the symmetric adjacency matrix $\mathbf{A}$, is symmetric, positive semidefinite, and has real eigenvalues in the range $[0,2]$.
Proof sketch; full proof in Appendix B.
This follows from spectral graph theory [6], as $\mathbf{A}$ is constructed to be symmetric and non-negative, and Eq. 6 defines the standard normalized Laplacian. ∎
3.3 Spectral Filtering Framework
To model user preferences at different levels of specificity, we perform spectral filtering on the Laplacian graph $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{T}$ where $\boldsymbol{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{m+n})$ contains eigenvalues in ascending order $0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{m+n}\leq 2$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{m+n}]$ contains the corresponding eigenvectors. In practice, for computational efficiency, we perform a truncated eigendecomposition focusing on $r\ll(m+n)$ which includes the low- and mid-frequency bands relevant to our filters. The lowpass filter $g_{\text{LP}}(\lambda)$ and bandpass filter $g_{\text{BP}}(\lambda)$ operate on the corresponding $r$ eigenpairs by selecting and amplifying the mid-frequency signals present within this subspace.
Proposition 3 (Laplacian Properties).
The eigenvalues of $\mathbf{L}$ lie in $[0,2]$, with the number of zero eigenvalues equal to the number of connected components in the graph.
Bandpass Filtering:
To capture mid-frequency user-specific behavior, we design a bandpass filter to the graph spectrum using a Gaussian kernel:
$g_{\text{BP}}(\lambda)=\exp\left(-\frac{(\bar{\lambda}-c)^{2}}{w}\right),\quad\bar{\lambda}=\frac{\lambda-\lambda_{\min}}{\lambda_{\max}-\lambda_{\min}}$ (7)
where $c$ and $w$ control the center and width of the frequency band. This form is motivated by our empirical analysis (Appendix G), which indicates that spectral coefficients for intermediate eigenvalues often show higher magnitude for user-specific preferences compared to global trends. Let $\mathbf{G}_{\text{BP}}=\text{diag}(g_{\text{BP}}(\lambda_{1}),\ldots,g_{\text{BP}}(\lambda_{r}))$ denote the spectral bandpass filter. We filter the item-degree-normalized user-item interaction signal using the band-pass operator:
$\mathbf{F}_{BP}=\mathbf{X}\mathbf{D}_{I}^{-1/2}\mathbf{U}\mathbf{G}_{BP}\mathbf{U}^{T}\mathbf{D}_{I}^{-1/2}$ (8)
Lowpass Filtering:
To preserve global structure and smooth preferences, we compute a complementary low-pass component augmented with user-user similarity. To enhance global structure, we augment user-item signals with user-user similarities. We define $\mathbf{X}_{b}=[\mathbf{C}_{U},\mathbf{X}]$, where $\mathbf{C}_{U}=\tilde{\mathbf{X}}_{U}\tilde{\mathbf{X}}_{U}^{T}$. The low-pass component is then computed as:
$\mathbf{F}_{LP}=\mathbf{X}_{b}\mathbf{D}_{b}^{-1/2}\mathbf{U}\mathbf{U}^{T}\mathbf{D}_{b}^{1/2}$ (9)
where $\mathbf{D}_{b}=\text{diag}(\mathbf{X}_{b}^{T}\mathbf{1}_{m})$. We note that $\mathbf{C}_{U}$ is not part of the Laplacian but is introduced to incorporate global user similarity in the lowpass representation. As $\mathbf{X}_{b}$ contains both user-user and user-item components, we extract only the item-related part after filtering by:
$\mathbf{F}_{LP}=\mathbf{F}_{LP}[:,m:]$ (10)
In our framework, mid-frequency components correspond to more user-specific variations, while low frequencies capture smoother, population-level preferences. We compute item scores by fusing mid-frequency personalization with low-frequency global trends. Specifically, we compute:
$\mathbf{Y}=\phi\cdot\mathbf{F}_{\text{BP}}+(1-\phi)\cdot\mathbf{F}_{\text{LP}},\quad\phi\in[0,1]$ (11)
[…]
4 Experiments
We evaluate GSPRec to answer: (1) How does sequential graph enrichment affect performance? (2) What is the impact of frequency-aware filtering? (3) How does GSPRec compare to state-of-the-art recommenders?
4.1 Setup
Data:
We conduct experiments on four benchmark datasets spanning two domains: (1) ML100K, ML1M [13] and Netflix [1] from the movie recommendation domain; (2) Amazon Beauty [32] from the e-commerce product domain. These datasets were selected to maintain comparability with recent spectral methods [52, 53] that established them as standard benchmarks for this class of approaches. All datasets are divided into training set and test set with the ratio of 8:2, and we take 10% of the training set as validation set for hyperparameter tuning. Table 3 shows the statistics of these datasets.
Baselines:
We compare GSPRec with: (i) Popularity; (ii) GCN-based methods: LightGCN [14], LR-GCCF [4], IMP-GCN [23], SimpleX [28], UltraGCN [29]; and (iii) GSP-based methods: GF-CF [42], PGSP [25], FaGSP [52], HiGSP [53]. A detailed description of these baselines is provided in Appendix E. Unlike existing GSP methods that use only collaborative signals and low-pass filtering, GSPRec encodes temporal transitions into the graph topology and employs a dual-filter approach targeting both low and bandpass filter components.
Implementation Details:
For graph construction, we set the diffusion depth $d=2$ and the decay $\alpha=0.4$ in all datasets to capture immediate and second-order transitions while attenuating longer-range influences. Eigendecomposition dimensionality $r$ varies by dataset size. Filter parameters $c$, $w$, and fusion weight $\phi$ are tuned per dataset as shown in Table 3. Implementation details are in Appendix F.
Table 2: Statistics of evaluation datasets.
Dataset #Users #Items #Inter. Density
ML100K 943 1,682 100,000 6.30%
ML1M 6,040 3,706 1,000,209 4.47%
Netflix 20,000 17,720 5,678,654 1.60%
Beauty 22,363 12,101 198,502 0.07%
Table 3: Filter parameters across datasets.
Dataset $\phi$ $c$ $w$ $r$
ML100K 0.5 0.2 0.1 32
ML1M 0.3 0.8 0.1 128
Netflix 0.3 0.4 0.3 256
Beauty 0.5 0.8 0.3 512
Metrics.
Following prior work [14, 42, 52, 53, 25], we use Normalized Discounted Cumulative Gain (NDCG@$k$) and Mean Reciprocal Rank (MRR@$k$) at $k\in\{5,10,20\}$. We apply spectral filtering to obtain preference scores, exclude previously interacted items, and rank the top-$k$ items.
4.2 Overall Performance Comparison
Table 4: Performance comparison on four public datasets. The best performance is denoted in bold. “M@K” and “N@K” are short for “MRR@K” and “NDCG@K”, respectively. “Improv.” denotes the percentage improvement of GSPRec compared to the strongest baseline method.
Datasets Metric GCN-based Methods GSP-based Methods Improv.
Popularity LightGCN LR-GCCF IMP-GCN SimpleX UltraGCN GF-CF PGSP FaGSP HiGSP GSPRec
ML100K N@5 0.5672 0.7034 0.5587 0.6764 0.6995 0.6786 0.6875 0.6851 0.7106 0.7166 0.7543 +5.26%
M@5 0.5151 0.6362 0.4943 0.6006 0.6397 0.6094 0.6234 0.6149 0.6518 0.6629 0.7358 +11.00%
N@10 0.5857 0.6771 0.5603 0.6605 0.6866 0.6688 0.6843 0.6722 0.7092 0.7021 0.7572 +6.77%
M@10 0.5296 0.5877 0.4616 0.5690 0.6064 0.5749 0.6010 0.5767 0.6384 0.6380 0.7443 +16.59%
N@20 0.5867 0.6618 0.5555 0.6605 0.6543 0.6486 0.6621 0.6495 0.6989 0.6724 0.7465 +6.81%
M@20 0.5315 0.5652 0.4411 0.5690 0.5411 0.5289 0.5593 0.5496 0.5901 0.5782 0.7457 +26.37%
ML1M N@5 0.0958 0.5845 0.3540 0.5583 0.5934 0.5739 0.5935 0.5963 0.6112 0.6062 0.6237 +2.05%
M@5 0.1943 0.5160 0.2997 0.4934 0.5244 0.5087 0.5254 0.5313 0.5430 0.5386 0.5715 +5.25%
N@10 0.0960 0.5873 0.3787 0.5594 0.5921 0.5773 0.5897 0.5923 0.6082 0.6042 0.6431 +5.74%
M@10 0.2180 0.5010 0.2946 0.4685 0.5051 0.4886 0.4996 0.5063 0.5229 0.5148 0.5837 +11.63%
N@20 0.1035 0.5692 0.3980 0.5416 0.5687 0.5600 0.5678 0.5687 0.5808 0.5769 0.6431 +10.73%
M@20 0.2309 0.4644 0.2924 0.4292 0.4625 0.4500 0.4557 0.4646 0.4748 0.4620 0.5871 +23.65%
Netflix N@5 0.5105 0.6822 0.6023 0.6328 0.6603 0.5680 0.7005 0.6811 0.7162 0.7162 0.7201 +0.54%
M@5 0.4505 0.6147 0.5323 0.5528 0.5902 0.4941 0.6360 0.6127 0.6524 0.6532 0.6680 +2.27%
N@10 0.5488 0.6814 0.6134 0.6307 0.6694 0.5746 0.6928 0.6756 0.7079 0.7062 0.7270 +2.70%
M@10 0.4712 0.5974 0.5234 0.5386 0.5839 0.4708 0.6126 0.5905 0.6293 0.6275 0.6762 +7.45%
N@20 0.5604 0.6642 0.6043 0.6152 0.6572 0.5580 0.6689 0.6543 0.6814 0.6800 0.7213 +5.86%
M@20 0.4773 0.5636 0.4963 0.4995 0.5555 0.4309 0.5691 0.5502 0.5827 0.5800 0.6777 +16.30%
Beauty N@5 0.0149 0.0668 0.0533 0.0570 0.0682 0.0599 0.0659 0.0674 0.0712 0.0718 0.0733 +2.09%
M@5 0.0118 0.0530 0.0423 0.0444 0.0544 0.0476 0.0523 0.0538 0.0564 0.0574 0.0619 +7.84%
N@10 0.0200 0.0763 0.0635 0.0667 0.0761 0.0649 0.0751 0.0765 0.0799 0.0771 0.0879 +10.01%
M@10 0.0139 0.0519 0.0443 0.0451 0.0517 0.0441 0.0517 0.0525 0.0549 0.0531 0.0680 +23.86%
N@20 0.0256 0.0829 0.0678 0.0729 0.0819 0.0689 0.0819 0.0828 0.0846 0.0799 0.1013 +19.74%
M@20 0.0154 0.0466 0.0388 0.0409 0.0461 0.0393 0.0469 0.0481 0.0485 0.0456 0.0718 +48.04%
Table 4 presents performance comparisons. GSPRec consistently outperforms all baselines across all metrics and datasets. Three Key observations: (1) consistent NDCG@$k$/MRR@$k$ improvements show stronger ranking quality, (2) larger gains on sparse datasets confirm robustness, and (3) improvements over recent GSP-based methods validate the benefit of leveraging both sequential and spectral signals.
4.3 Parameter Study
Sequential-Aware Graph:
To isolate the contribution of sequential graph, we compare our full model with four variants shown in Table 6: (1) GSPRec-NB: no band-pass filter, (2) GSPRec-NL: no low-pass filter, (3) GSPRec-NS: no sequential diffusion, (4) GSPRec-SE: sequence edges only (no diffusion). The non-sequential variant (GSPRec-NS) shows significant performance drops, confirming the value of sequential information. The direct-sequence variant (GSPRec-SE) performs slightly worse than the full model, demonstrating that higher-order transitions captured by multi-hop diffusion provide additional benefits.
Figure 4: Frequency response curves of different filter configurations all datasets. The low-pass filter (blue) preserves signals in the lower 30% of the spectrum. The band-pass filter (green) selectively amplifies mid-frequency components associated with personalized patterns. The combined filter response (purple) illustrates how dual-filter captures complementary signals across the spectrum.
Table 5: Ablation study on ML1M dataset.
Variant N@10 $\pm$ SE M@10 $\pm$ SE
GSPRec-NB 0.5769 0.0082 0.5148 0.0093
GSPRec-NL 0.6042 0.0079 0.5229 0.0088
GSPRec-NS 0.6274 0.0039 0.5704 0.0050
GSPRec-SE 0.6416 0.0037 0.5831 0.0049
GSPRec 0.6431 0.0074 0.5837 0.0081
Table 6: Filtering comparison on ML1M dataset.
Components N@10 $\pm$ SE M@10 $\pm$ SE
$\mathbf{F}_{\text{LP}}$ only 0.5769 0.0082 0.5148 0.0093
$\mathbf{F}_{\text{BP}}$ only 0.6042 0.0079 0.5229 0.0088
$\mathbf{F}_{\text{LP}}$+$\mathbf{F}_{\text{BP}}$ ($\phi$=0.3) 0.6431 0.0074 0.5837 0.0081
$\mathbf{F}_{\text{LP}}$+$\mathbf{F}_{\text{BP}}$ ($\phi$=0.5) 0.6327 0.0076 0.5714 0.0085
$\mathbf{F}_{\text{LP}}$+$\mathbf{F}_{\text{BP}}$ ($\phi$=0.7) 0.6281 0.0078 0.5682 0.0087
Spectral Filtering Performance:
Table 3 shows optimal filter settings across datasets: (1) larger datasets require higher eigendecomposition dimensions $r$; (2) sparser datasets benefit from wider bandpass filters; and (3) fusion weight $\phi$ correlates with dataset density. Figure 4 illustrates filter frequency responses spanning both global and user-specific bands. As shown in Table 6, the dual-filter approach consistently outperforms single-filter alternatives, with best performance at $\phi$=0.3 on ML1M. Notably, the bandpass filter alone outperforms the lowpass filter alone by 4.7% in N@10 and 1.6% in M@10, highlighting the importance of mid-frequency components for personalization. Ablation results in Table 6 further support this, with GSPRec-NB (no band-pass) showing the largest performance drops.
4.4 Parameter Sensitivity Analysis
Figure 5: Effect of bandpass center position ($c$) on NDCG@10 across datasets. Each line represents a different filter bandwidth ($w$).
Figure 6: Effect of bandpass width ($w$) on NDCG@10 across datasets. Each line represents a different center position ($c$). Optimal width depends on center position and dataset characteristics.
Figures 5 and 6 show how performance varies with band-pass filter parameters. The center position ($c$) exhibits dataset-specific optima, reflecting varying spectral distributions across recommendation domains. Filter width ($w$) interacts with center position. The fusion weight $\phi$ also shows dataset-dependent patterns: denser datasets favor global signals (lower $\phi$), while sparser datasets benefit from stronger mid-frequency personalization (higher $\phi$).
5 Conclusion
We introduced GSPRec, a graph–spectral framework that jointly models collaborative and sequential signals via a unified Laplacian. Our method augments the user–item graph with multi-hop, exponentially decayed item transitions and applies dual spectral filters. Theoretically, we ensure spectral convergence and Laplacian validity. Empirically, we demonstrate consistent improvements across four benchmarks over state-of-the-art GSP methods. Ablation studies confirm the necessity of both sequential graph encoding and frequency-aware filtering. Future work includes extending to directed graphs, learning adaptive spectral filters, and applying to diverse recommendation contexts.
6 Broader Impact
This work improves recommender systems by integrating sequential behavior with graph spectral analysis, benefiting users across e-commerce, media, and education domains. However, spectral methods may amplify biases in historical data, potentially reinforcing popularity bias or increasing computational overhead. Future work should explore fairness-aware spectral filters, robustness evaluations, and privacy-preserving techniques. We encourage practitioners to consider these tradeoffs when deploying these systems in real-world settings.
Appendices
This supplementary material provides detailed information that complements our main paper. It includes formal notation definitions, theoretical analysis, algorithmic details, computational complexity analysis, and experimental information to support reproducibility.
Appendix A Table of Notations
We summarize the key notations used throughout this paper in Table 7.
Table 7: Summary of Notations
Notation Description
Input Data Representations
$\mathcal{U},\mathcal{I}$ Set of users and items, respectively
$\mathcal{D}$ Historical interactions as $(u,i,t)$ triplets
$\mathbf{X}\in\{0,1\}^{m\times n}$ Binary user-item interaction matrix
$\mathcal{S}_{u}$ Time-ordered sequence of items for user $u$
Graph Construction
$\mathbf{S}$ Initial directed item transition matrix from $\mathcal{S}_{u}$
$\mathbf{S}^{\prime}$ Symmetric item connectivity graph derived from $\mathbf{S}$
$\alpha,d$ Diffusion decay factor and depth
$\mathbf{S}^{(d)}$ Multi-hop diffusion matrix from $\mathbf{S}^{\prime}$ up to depth $d$
$\tilde{\mathbf{S}}$ Normalized diffusion matrix derived from $\mathbf{S}^{(d)}$
$\mathbf{A}$ Unified user-item adjacency matrix
$\mathbf{L}$ Normalized Laplacian matrix
Spectral Filtering
$\mathbf{U},\boldsymbol{\Lambda}$ Matrices of eigenvectors and eigenvalues of $\mathbf{L}$, respectively
$r$ Number of eigenvectors in truncated decomposition
$g_{\text{BP}}(\lambda)$ Gaussian bandpass filter kernel
$c,w$ Center and width of bandpass filter
$\mathbf{F}_{\text{BP}}$ Bandpass filtered user-item signal
$\mathbf{F}_{\text{LP}}$ Lowpass filtered user-item signal
$\phi$ Fusion parameter balancing filter components
$\mathbf{Y}$ Final prediction matrix
Appendix B Algorithmic Details
B.1 GSPRec Algorithm
The full end-to-end procedure of our method is outlined in Algorithm 1. It summarizes the construction of the sequential graph encoding, Laplacian computation, spectral filtering, and final score generation.
Input: Interaction set $\mathcal{D}=\{(u,i,t)\}$, diffusion order $d$, decay factor $\alpha$, filter parameters $(c,w)$, fusion weight $\phi$
Output: Predicted score matrix $\mathbf{Y}\in\mathbb{R}^{m\times n}$
// Graph Construction
$\mathbf{X}\leftarrow$ Binary user-item matrix from $\mathcal{D}$
$\mathcal{S}_{u}\leftarrow$ Chronologically sorted item sequences per user
$\mathbf{S}\leftarrow$ Initial directed item transition matrix from $\mathcal{S}_{u}$ (where $S_{ij}=1$ if $i\to j$)
$\mathbf{S}^{\prime}\leftarrow\text{Symmetrize}(\mathbf{S})$ // $S^{\prime}_{ij}=1$ if $S_{ij}=1\text{ or }S_{ji}=1$
$\mathbf{S}^{(d)}\leftarrow\sum_{k=1}^{d}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$ // Exponential diffusion on symmetric $\mathbf{S}^{\prime}$
$\tilde{\mathbf{S}}\leftarrow\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2}$ // where $\mathbf{D}_{S}=\text{diag}(\mathbf{S}^{(d)}\mathbf{1}_{n})$
$\mathbf{A}\leftarrow\begin{bmatrix}\mathbf{0}&\mathbf{X}\\ \mathbf{X}^{\top}&\tilde{\mathbf{S}}\end{bmatrix}$
$\mathbf{D}\leftarrow\text{diag}(\mathbf{A}\mathbf{1})$
$\mathbf{L}\leftarrow\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
$[\mathbf{U},\mathbf{\Lambda}]\leftarrow$ Eigendecomposition of $\mathbf{L}$
// Frequency-Aware Spectral Filtering
$\bar{\boldsymbol{\lambda}}\leftarrow$ Diagonal entries of $\mathbf{\Lambda}$
$\mathbf{G}_{\text{BP}}\leftarrow\text{diag}\left(\exp\left(-\frac{(\bar{\boldsymbol{\lambda}}-c)^{2}}{w}\right)\right)$ // Gaussian band-pass
$\mathbf{F}_{\text{BP}}\leftarrow\mathbf{X}\mathbf{D}_{I}^{-1/2}\mathbf{U}\mathbf{G}_{\text{BP}}\mathbf{U}^{\top}\mathbf{D}_{I}^{-1/2}$
// Low-Pass Filtering
$\tilde{\mathbf{X}}_{U}\leftarrow\mathbf{D}_{U}^{-1/2}\mathbf{X}$ // where $\mathbf{D}_{U}=\text{diag}(\mathbf{X1}_{n})$, $\mathbf{D}_{I}=\text{diag}(\mathbf{X}^{T}\mathbf{1}_{m})$
$\mathbf{C}_{U}\leftarrow\tilde{\mathbf{X}}_{U}\tilde{\mathbf{X}}_{U}^{\top}$
$\mathbf{X}_{b}\leftarrow[\mathbf{C}_{U},\mathbf{X}]$
$\mathbf{D}_{b}\leftarrow\text{diag}(\mathbf{X}_{b}^{\top}\mathbf{1})$
$\mathbf{F}_{\text{LP}}\leftarrow\mathbf{X}_{b}\mathbf{D}_{b}^{-1/2}\mathbf{U}\mathbf{U}^{\top}\mathbf{D}_{b}^{1/2}$
$\mathbf{F}_{\text{LP}}\leftarrow\mathbf{F}_{\text{LP}}[:,m{:}]$ // Select only item columns
// Fusion and Output
$\mathbf{Y}\leftarrow\phi\cdot\mathbf{F}_{\text{BP}}+(1-\phi)\cdot\mathbf{F}_{\text{LP}}$
return $\mathbf{Y}$
Algorithm 1 GSPRec
B.2 Proof of Proposition 2: Laplacian Validity
Proof.
We first establish properties of the adjacency matrix $\mathbf{A}\in\mathbb{R}^{(m+n)\times(m+n)}$, defined as:
$\mathbf{A}=\begin{pmatrix}\mathbf{0}_{m\times m}&\mathbf{X}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}\end{pmatrix}$ (12)
where $\mathbf{X}\in\{0,1\}^{m\times n}$ (the binary user-item interaction matrix) and $\tilde{\mathbf{S}}=\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2}\in\mathbb{R}^{n\times n}$ (the normalized sequence-derived item-item graph, with $\mathbf{S}^{(d)}$ being the diffused matrix from Eq. 2).
Lemma 4.
The adjacency matrix $\mathbf{A}$ is symmetric and non-negative.
Proof of Lemma 4.
To prove symmetry of $\mathbf{A}$, we note:
$\mathbf{A}^{T}$ $=\begin{pmatrix}\mathbf{0}_{m\times m}^{T}&(\mathbf{X}^{T})^{T}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}^{T}\end{pmatrix}$ (13)
$=\begin{pmatrix}\mathbf{0}_{m\times m}&\mathbf{X}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}\end{pmatrix}=\mathbf{A}$ (14)
This equality holds because $\mathbf{0}_{m\times m}$ is symmetric, $(\mathbf{X}^{T})^{T}=\mathbf{X}$, and $\tilde{\mathbf{S}}^{T}=\tilde{\mathbf{S}}$. The matrix $\tilde{\mathbf{S}}$ is symmetric by construction: the base matrix $\mathbf{S}^{\prime}$ used in the diffusion is symmetric, thus $\mathbf{S}^{(d)}$ (from Eq. 2) is symmetric, and the symmetric normalization (Eq. 3) preserves this symmetry.
For non-negativity of $\mathbf{A}$, we show all its elements are non-negative:
  • $\mathbf{X}\in\{0,1\}^{m\times n}$ by definition, so its elements are $\geq 0$. Thus, $\mathbf{X}\geq 0$ and $\mathbf{X}^{T}\geq 0$.
  • For $\tilde{\mathbf{S}}$: The initial directed transition matrix $\mathbf{S}$ has $s_{ij}\in\{0,1\}$, so $\mathbf{S}\geq 0$. The symmetric base graph $\mathbf{S}^{\prime}$ is constructed from $\mathbf{S}$ such that its entries $s^{\prime}_{ij}\in\{0,1\}$, so $\mathbf{S}^{\prime}\geq 0$. Given $\mathbf{S}^{\prime}\geq 0$ and the decay factor $\alpha\in(0,1)>0$, each power $(\mathbf{S}^{\prime})^{k}$ in the diffusion sum will have non-negative entries. Therefore, $\mathbf{S}^{(d)}=\sum_{k=1}^{d}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$ (Eq. 2) will also have non-negative entries, i.e., $\mathbf{S}^{(d)}\geq 0$. The degree matrix $\mathbf{D}_{S}=\text{diag}(\mathbf{S}^{(d)}\mathbf{1}_{n})$ will have non-negative diagonal entries. For non-isolated nodes in the graph represented by $\mathbf{S}^{(d)}$, these entries will be positive, making $\mathbf{D}_{S}^{-1/2}$ well-defined with non-negative real entries. Hence, $\tilde{\mathbf{S}}=\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2}\geq 0$.
Therefore, $\mathbf{A}$ is both symmetric and non-negative. ∎
Given that $\mathbf{A}$ is symmetric and non-negative (by Lemma 4), and its degree matrix $\mathbf{D}=\text{diag}(\mathbf{A}\mathbf{1})$, the normalized Laplacian is $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ (Eq. 6). We apply Theorem 5:
Theorem 5 (Properties of Normalized Laplacian (Chung, 1997 [6])).
For a symmetric, non-negative adjacency matrix $\mathbf{A}$ with degree matrix $\mathbf{D}$, the normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ satisfies: (1) $\mathbf{L}$ is real and symmetric. (2) $\mathbf{L}$ is positive semi-definite. (3) All eigenvalues of $\mathbf{L}$ lie in $[0,2]$.
Applying Theorem 5 directly to our $\mathbf{A}$ and $\mathbf{L}$ concludes that $\mathbf{L}$ is symmetric, positive semi-definite, and has eigenvalues in $[0,2]$. ∎
Appendix C Theoretical Analysis of Spectral Bands
The concentration of user-specific patterns in mid-frequency components can be understood through the lens of graph signal smoothness. Low-frequency components correspond to signals that are smooth across the graph, capturing global popularity trends that vary slowly across the user-item structure. High-frequency components often represent noise or sparse outlier preferences.
Mid-frequency components represent a balance where signals vary moderately across local graph neighborhoods - precisely the pattern exhibited by personalized preferences that differ between user communities but show consistency within them. This aligns with previous findings in spectral clustering [47] and network community detection [31], where mid-frequency eigenvectors have been shown to encode community structure.
We can formalize this intuition through the concept of total variation of a signal [43]. Given a signal $\mathbf{f}$ on a graph with Laplacian $\mathbf{L}$, the quadratic form $\mathbf{f}^{T}\mathbf{L}\mathbf{f}$ measures how smoothly the signal varies across the graph. For the eigenvectors $\mathbf{u}_{i}$ of $\mathbf{L}$, we have $\mathbf{u}_{i}^{T}\mathbf{L}\mathbf{u}_{i}=\lambda_{i}$, where $\lambda_{i}$ is the corresponding eigenvalue.
User preference signals can be decomposed into these eigenvectors:
$\mathbf{f}=\sum_{i=1}^{m+n}\alpha_{i}\mathbf{u}_{i}$ (15)
In our empirical analysis across datasets, we consistently observe that coefficients $\alpha_{i}$ corresponding to mid-frequency components ($0.80.3<\lambda_{i}<$ 0.8) have significantly higher magnitude for user-specific preference patterns compared to global trends. This provides empirical validation of our theoretical understanding of spectral representation in recommendation graphs.
Appendix D Computational Complexity Analysis
We analyze the computational complexity of GSPRec with $m$ users, $n$ items, $|\mathcal{D}|$ interactions, diffusion depth $d$, and $r$ eigenvectors.
D.1 Time Complexity
The time complexity of key components is:
  • Graph Construction: $O(|\mathcal{D}|+d\cdot\text{nnz}(\mathbf{S}^{\prime})\cdot n)$ for building and diffusing the sequential transition matrix.
  • Eigendecomposition: $O(r\cdot(m+n)^{2})$ for computing the $r$ smallest eigenvectors of the Laplacian.
  • Spectral Filtering: $O(|\mathcal{D}|\cdot r+(m+n)\cdot r^{2})$ for applying band-pass and low-pass filters.
The overall time complexity is dominated by the eigendecomposition step: $O(r\cdot(m+n)^{2})$. For sparse interaction data and $r\ll(m+n)$, the algorithm remains computationally tractable.
D.2 Space Complexity
The space requirements are:
  • Input Data: $O(|\mathcal{D}|+\text{nnz}(\mathbf{S}^{(d)}))$ for storing user-item interactions and sequential transitions.
  • Spectral Representation: $O(r\cdot(m+n))$ for storing eigenvectors.
  • Output: $O(m\cdot n)$ for the filtered results and final predictions.
The overall space complexity is $O(|\mathcal{D}|+\text{nnz}(\mathbf{S}^{(d)})+r\cdot(m+n)+m\cdot n)$, which is manageable for sparse real-world recommendation datasets where $|\mathcal{D}|\ll m\cdot n$ and $\text{nnz}(\mathbf{S}^{(d)})\ll n^{2}$.
In practice, the algorithm can be further optimized through sparse matrix operations, approximate eigendecomposition methods, and parallelization of filtering operations.
Appendix E Baseline Details
We compare GSPRec with a comprehensive set of baselines representing both GCN-based and GSP-based CF methods:
GCN-based CF Methods:
  • LightGCN [14] simplifies GCN for recommendation by removing feature transformation and non-linear activation, retaining only essential neighborhood aggregation.
  • LR-GCCF [4] introduces residual connections to graph convolutional networks for improving recommendation performance.
  • IMP-GCN [23] propagates information in user sub-graphs to reduce the impact of noise and alleviate the over-smoothing problem.
  • SimpleX [28] focuses on optimizing loss function and negative sampling strategies.
  • UltraGCN [29] approximates infinite-layer graph convolutions via constraint loss and flexible edge weight assignment.
GSP-based CF Methods:
  • GF-CF [42] integrates linear and ideal low-pass filters to model user preferences.
  • PGSP [25] uses a mixed-frequency low-pass filter over personalized graph signals.
  • FaGSP [52] employs Cascaded and Parallel Filter Modules for user preference modeling.
  • HiGSP [53] adopts cluster-wise and globally-aware filters to recognize unique and general interaction patterns.
For comprehensive evaluation, we report results for LR-GCCF, IMP-GCN, SimpleX, UltraGCN, and GF-CF as presented in [53, 52], which used identical dataset preprocessing and evaluation protocols.
Appendix F Reproducibility
To improve reproducibility, we provide the complete algorithm of GSPRec in Algorithm 1, which details the graph construction, spectral decomposition, and dual filtering processes. The optimal hyperparameters for all experimental datasets are summarized in Table 3. For implementation, we process datasets following the procedure described in Section 4.1, which involves splitting datasets in an 8:2 ratio for training/testing, with 10% of the training set used for validation. When implementing GSPRec, the key parameter settings are:
  • Diffusion depth $d=2$ and decay factor $\alpha=0.4$ for sequential graph construction
  • Dataset-specific eigenvector count $r$, band-pass center $c$, width $w$, and fusion weight $\phi$ as specified in Table 3
For evaluation, we follow the protocol detailed in Section 4.1, using NDCG and MRR metrics at $k\in\{5,10,20\}$ after excluding previously interacted items.
F.1 Computational Resources
Experiments were conducted on a system with AMD EPYC processors and NVIDIA RTX A6000 GPUs. The computational requirements varied with dataset size and the number of eigenvectors used. The eigendecomposition step is the primary computational bottleneck as identified in our complexity analysis (Appendix D). For benchmark datasets used in our experiments, the runtime and memory requirements were reasonable, allowing for comprehensive evaluation on standard research hardware. Our experimental protocol, including parameter tuning and ablation studies, involved multiple runs for each configuration to ensure statistical significance.
F.2 Scaling to Large-Scale Systems
For industrial-scale applications with millions of users and items, we propose three optimization strategies:
F.2.1 Approximate Eigendecomposition
The dominant computational bottleneck can be addressed using the Nyström method [8] or randomized SVD approaches [11], reducing complexity from $O(r(m+n)^{2})$ to $O(r^{2}(m+n))$ (or better, depending on specific variant, e.g., linear in $(m+n)$ for matrix-vector products), with minimal performance impact.
F.2.2 Graph Sparsification
For large-scale graphs, we can apply spectral sparsification techniques [44] that preserve the essential spectral properties while reducing edge density, accelerating both graph construction and eigendecomposition.
F.2.3 Incremental Updates
In production environments, the full spectral decomposition need not be recomputed for each recommendation. We propose an incremental update scheme where the spectral components are updated periodically while incremental preference updates can be applied in real-time.
Appendix G Mathematical Analysis of Bandpass Filtering
We provide mathematical justification for our bandpass filtering approach and its relevance to recommendation systems. Our Gaussian kernel on the graph Laplacian eigenspectrum is defined as in Equation 7.
$g(\lambda)=\exp\left(-\frac{(\bar{\lambda}-c)^{2}}{w}\right),\quad\bar{\lambda}=\frac{\lambda-\lambda_{\min}}{\lambda_{\max}-\lambda_{\min}}$
This satisfies the formal definition of a bandpass filter: it achieves maximum response when $\bar{\lambda}=c$ (which corresponds to $\lambda_{c}=c(\lambda_{\max}-\lambda_{\min})+\lambda_{\min}$ in the original eigenvalue domain) and attenuates components with eigenvalues far from $\lambda_{c}$. For our normalized Laplacian with original eigenvalues $\lambda\in[0,2]$ (so $\lambda_{\min}\approx 0,\lambda_{\max}\approx 2$), setting $c\in(0,1)$ (the center in the normalized $\bar{\lambda}\in[0,1]$ domain) ensures the filter passes a band of mid-range eigenvalues while attenuating both low and high extremes. When applied to user preference signals decomposed as $\mathbf{x}=\sum_{i=1}^{m+n}\alpha_{i}\mathbf{u}_{i}$, the filtered output $\mathbf{y}=\mathbf{U}\text{diag}(g(\boldsymbol{\Lambda}))\mathbf{U}^{T}\mathbf{x}$ selectively preserves components $\alpha_{i}\mathbf{u}_{i}$ where the eigenvalue $\lambda_{i}$ (corresponding to $\mathbf{u}_{i}$) results in a high $g(\lambda_{i})$ value.
Our empirical analysis across recommendation datasets reveals that coefficients $\alpha_{i}$ corresponding to intermediate eigenvalues ($0.80.3<\lambda_{i}<$ 0.8, assuming $\lambda_{i}$ are original eigenvalues in $[0,2]$) have higher magnitude for user-specific preference patterns compared to global trends. This confirms that mid-frequency components encode valuable personalization signals that traditional low-pass approaches would suppress.
Table 6 provides quantitative evidence of this phenomenon: while low-pass filtering captures popularity trends, our bandpass approach more effectively preserves the user-specific patterns essential for accurate personalized recommendations.