### v2

# GSPRec: Temporal-Aware Graph Spectral Filtering for Recommendation

Abstract
Graph-based recommendation systems are effective at modeling collaborative patterns but often suffer from two limitations: overreliance on low-pass filtering, which suppresses user-specific signals, and omission of sequential dynamics in graph construction. We introduce GSPRec, a graph spectral model that integrates temporal transitions through sequentially-informed graph construction and applies frequency-aware filtering in the spectral domain. GSPRec encodes item transitions via multi-hop diffusion to enable the use of symmetric Laplacians for spectral processing. To capture user preferences, we design a dual-filtering mechanism: a Gaussian bandpass filter to extract mid-frequency, user-level patterns, and a low-pass filter to retain global trends. Extensive experiments on four public datasets show that GSPRec consistently outperforms baselines, with an average improvement of 6.77% in NDCG@10. Ablation studies show the complementary benefits of both sequential graph augmentation and bandpass filtering.
1 Introduction
Graph-based recommender systems are constrained by two core problems: using low-pass spectral filtering often suppresses distinct user signals in favor of broad popularity patterns, and the order of user interactions is typically neglected in their graph structures. Therefore, recommender systems should leverage both collaborative patterns and the rich temporal signals embedded within these interaction sequences Koren (2009). Graph-based Collaborative Filtering (CF) methods He et al. (2020); Shen et al. (2021); Wang et al. (2019) have achieved strong performance, but they face two limitations. First, models such as LightGCN He et al. (2020), which rely on linear embedding propagation, act as low-pass filters Shen et al. (2021) by smoothing signals across the user-item graph. This encourages broad popularity effects (e.g., everyone buys laptops) but suppresses subgroup-specific patterns (e.g., gamers buy laptops with mice, artists buy laptops with tabletsShen et al. (2021); Xia et al. (2025). Second, while some GCN-based recommenders incorporate temporal data, they do so via separate sequential modules Ma et al. (2020); Hsu & Li (2021) or through methods that do not integrate order into the graph structure for spectral analysis, which means they fail to preserve sequential information essential for distinguishing user preferences from popularity patterns.
The filtering behavior inherent in many GCNs motivates a closer look through the lens of graph signal processing (GSP), which generalizes classical signal processing to irregular graphs Ortega et al. (2018). A core technique in GSP is spectral filtering, which transforms graph signals using the eigenbasis of the graph Laplacian Kruzick & Moura (2017). While low-pass filters, widely adopted under the assumption of smoothness Shen et al. (2021); He et al. (2020), capture global popularity, they fail to retain the more distinctive user-specific variations Liu et al. (2023); Xia et al. (2025). Recent methods have explored amplifying high frequencies Guo et al. (2023), but risk magnifying noise. By contrast, mid-frequency components encode localized, task-relevant variations Dong et al. (2019) remains underexplored, motivating filters that explicitly target this band for personalization.
We introduce GSPRec, a novel GSP-based framework that achieves temporal awareness in spectral recommenders through two components: First, its sequentially-aware graph construction encodes multi-hop item transition patterns from user histories via diffusion into a symmetric graph structure to make temporal dynamics an intrinsic property for spectral analysis. Second, this graph empowers a frequency-aware dual-filtering by using Gaussian bandpass filter to extract mid-frequency user-specific preferences, complemented by a low-pass filter for popularity patterns.
In this work, we encode the interaction graph with item-to-item transitions derived from user interaction sequences. A multi-hop diffusion process with exponential decay captures sequential dependencies while preserving the benefits of symmetric graph structures. The symmetric structure guarantees real eigenvalues and orthogonal eigenvectors Chung (1997), which are important for stable spectral decomposition and well-defined filtering operations Ortega et al. (2018). Unlike prior work that extracts item-item similarity from collaborative information Xia et al. (2024), our approach aggregates explicit sequential signals into the graph topology [^1] We note that sequences are used for graph construction, not for sequential next-item prediction.. Figure 1 illustrates this shift from sequence-based to graph-based modeling.
Moreover, we design a Gaussian bandpass filter tailored to recommendation graphs. Informed by empirical spectral energy distributions Qin et al. (2024); Tremblay et al. (2018), the filter selectively captures mid-frequency components that are often associated with user-specific and sequential patterns positioned between low-frequency popularity signals and high-frequency noise. The Gaussian profile provides optimal localization in both spectral and vertex domains Shuman et al. (2013); Ortega et al. (2018), which enables smooth transitions across frequencies and reduces ringing artifacts commonly introduced by sharp filter cutoffs Hammond et al. (2011). We summarize our contributions as follows:
(a) Sequential pattern
(b) Bidirectional pattern
(c) Undirected graph
Figure 1: Conceptual shift from sequence-based to graph-based modeling. (a) Unidirectional transitions in RNNs Hidasi et al. (2015); Kang & McAuley (2018). (b) Bidirectional models like BERT4Rec Sun et al. (2019). (c) Our undirected item graph construction: solid lines show initial symmetric graph $\mathbf{S}^{\prime}$, dashed lines show higher-order relationships from multi-hop diffusion.
  • We propose a novel graph construction method that first derives a symmetric item connectivity graph from user sequences, then encodes higher-order sequential dependencies via multi-hop diffusion on this symmetric structure.
  • We propose a Gaussian bandpass filter that targets mid-frequency components.
  • We provide theoretical analysis establishing the stability and convergence of our sequential diffusion and filtering mechanisms.
  • We validate our framework on four public benchmarks, showing that integrating sequential transitions with mid-frequency filtering leads to consistent performance gains over baselines.
GSPRec achieves average gains of 6.77% in NDCG@10 and 14.88% in MRR@10 across four public benchmark datasets, consistently outperforming state-of-the-art baselines. These results highlight the effectiveness of our two core design choices: sequential graph encoding and spectral filtering.
2 Background & Related Work
Graph Signal Processing Background:
Graph Signal Processing (GSP) analyzes signals on graph structures Ortega et al. (2018); Shuman et al. (2013). A graph signal maps values to nodes $\mathcal{G}=(\mathcal{V},\mathcal{E})$ as $\mathbf{x}\in\mathbb{R}^{|\mathcal{V}|}$. The normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ decomposes into eigenvalues $\bm{\Lambda}$ and eigenvectors $\mathbf{U}$ Chung (1997), enabling spectral filtering:
$\mathbf{x}^{\prime}=\mathbf{U}\,g(\bm{\Lambda})\,\mathbf{U}^{T}\mathbf{x},\quad\text{where }g(\bm{\Lambda})=\mathrm{diag}(g(\lambda_{1}),\ldots,g(\lambda_{n}))$ (1)
Low-pass filters preserve global patterns, while band-pass filters capture mid-frequency variations related to user-specific behavior Ortega et al. (2018); Hammond et al. (2011).
Graph-Based Collaborative Filtering:
Graph-based collaborative filtering represents users and items as nodes in a bipartite graph, modeling interactions through various approaches He et al. (2020); Wang et al. (2019); Ying et al. (2018); Zheng et al. (2018). Recent works address over-smoothing Chen et al. (2020a); Shen et al. (2021), popularity bias Zhang et al. (2023), and efficiency Liang et al. (2024), while others incorporate temporal dynamics through sequential models or temporal graphs Kang & McAuley (2018); Sun et al. (2019); Zhang et al. (2024); Ou et al. (2025). Our approach encodes temporal locality as edge-level structure and applies spectral filtering without message passing.
GSP in Recommendation:
Prior work adapts GSP to recommender systems by treating interactions as graph signals. Most methods adopt low-pass filtering Huang et al. (2017); Liu et al. (2023); Xia et al. (2025; 2024), while recent work Guo et al. (2023) explores high-pass filtering, which may recover details but also amplify noise Shuman et al. (2013); Ortega et al. (2018). These approaches operate on static user-item graphs without sequential dynamics. In this work, we integrate sequential transitions into the graph before spectral analysis.
Table 1: Comparison with existing popular methods.
Method Task Spectral Filtering Temporal Info. Graph Structure Message Passing Mid-Freq
Message Passing GCNs:
NGCF Wang et al. (2019) General User-Item
LightGCN He et al. (2020) General User-Item
LR-GCCF Chen et al. (2020b) General User-Item
IMP-GCN Liu et al. (2021) General User-Item
Spectral Filtering Methods (GSP-based):
GF-CF Shen et al. (2021) General LP User-Item
JGCF Guo et al. (2023) General HP User-Item
PGSP Liu et al. (2023) General LP User-Item
HiGSP Xia et al. (2024) General LP User-Item
FaGSP Xia et al. (2025) General LP User-Item
Sequential Recommendation:
BERT4Rec Sun et al. (2019) Sequential Sequence
SASRec Kang & McAuley (2018) Sequential Sequence
GSPRec (Ours) General BP+LP UI+Seq
Table 1 summarizes existing approaches. Unlike GCN-based methods that rely on message passing, or GSP-based methods that apply low-pass (LP) or high-pass (HP) filtering to graphs, we join temporal graph construction (UI+Seq) with dual spectral filtering (BP+LP) to target mid-frequency components.
3 GSPRec
GSPRec consists of a sequential graph encoding procedure and a frequency-aware spectral filtering framework. These components jointly capture two complementary signal patterns in the spectral domain: (1) popularity patterns across the entire user-item graph, and (2) personalized patterns that distinguish individual preferences. Figure 2 summarizes the pipeline from user-item interactions to final recommendations via graph construction and spectral filtering.
Figure 2: Overview of GSPRec. Starting from user-item interaction sequences (left), we construct a unified graph joining user-item links and item-item transitions via multi-hop diffusion with exponential decay (Eq. 2). This graph is processed using two complementary spectral filters: a band-pass filter to extract mid-frequency, user-specific patterns, and a low-pass filter to capture popularity.
3.1 Problem Formulation
Let $\mathcal{U}=\{u_{1},\ldots,u_{m}\}$ denote users and $\mathcal{I}=\{i_{1},\ldots,i_{n}\}$ denote items. Historical interactions are triplets $\mathcal{D}=\{(u,i,t)\mid u\in\mathcal{U},i\in\mathcal{I},t\in\mathbb{R}^{+}\}$, yielding binary matrix $\mathbf{X}\in\{0,1\}^{m\times n}$ where $x_{ui}=1$ if user $u$ interacted with item $i$. For each user $u$, we define a time-ordered sequence $\mathcal{S}_{u}=[i_{1},i_{2},\ldots]$ where $(u,i_{j},t_{j})\in\mathcal{D}$ and $t_{1}<t_{2}<\cdots$. Our task is to predict future user-item interactions by jointly modeling collaborative patterns and sequential dynamics. We aim to predict unobserved entries in $\mathbf{X}$ using a unified spectral graph approach that leverages both user preference similarities and temporal item transitions.
3.2 Recommendation Graph Construction
Our goal is to construct a unified graph that integrates collaborative filtering signals with sequential transition patterns. Graph signal processing requires symmetric Laplacians to guarantee real eigenvalues, orthogonal eigenvectors, and stable spectral filtering Chung (1997); Ortega et al. (2018). Although user interaction histories are observed as timestamped sequences and thus appear directed, rigid order is not always reliable in recommendation domains Sun et al. (2019). We therefore transform sequences into a symmetric item-item graph component $\tilde{\mathbf{S}}$, a step that is theoretically required for spectral analysis. Multi-hop diffusion further preserves sequential proximity within this symmetric structure, allowing temporal information to remain embedded in the spectral representation.
Figure 1 illustrates our approach: while models in Figures 1(a) and 1(b) process sequences through uni/bidirectional mechanisms, we transform user interaction sequences into the undirected item graph in Figure 1(c). This enables joint analysis of collaborative and sequential signals with well-defined spectral properties Kruzick & Moura (2017). Direct use of asymmetric Laplacians Chung (2005) for directed graphs involves different theoretical formulation left for future work.
Initial Symmetric Sequence Graph ($\mathbf{S}^{\prime}$):
We define an initial directed transition matrix $\mathbf{S}\in\{0,1\}^{n\times n}$, where $s_{ij}=1$ if item $i$ directly precedes item $j$ in any user sequence. To adapt this for symmetric spectral methods, we form its undirected version, $\mathbf{S}^{\prime}\in\{0,1\}^{n\times n}$. In $\mathbf{S}^{\prime}$, an undirected edge $(i,j)$ exists if a transition $i\rightarrow j$ or $j\rightarrow i$ is in $\mathbf{S}$. This symmetrization is a standard practice when preparing directed graph for algorithms requiring undirected structures Von Luxburg (2007); Satuluri & Parthasarathy (2011). While simplifying transition directionality, this approach allows capturing broader item relatedness from sequences, as rigid order is not always optimal and wider context can be beneficial Sun et al. (2019). Multi-hop diffusion (Eq. 2) operates on $\mathbf{S}^{\prime}$ to encode sequence-derived item associations.
[…]
Multi-Hop Diffusion ($\mathbf{S}^{(d)}$):
To model broader item influence through multi-step paths, we apply multi-hop diffusion to $\mathbf{S}^{\prime}$:
$\mathbf{S}^{(d)}=\sum_{k=1}^{d}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$ (2)
where $\alpha\in(0,1)$ is an exponential decay factor that gives higher weight to shorter paths in $\mathbf{S}^{\prime}$, thus emphasizing local sequential relationships. $d$ is the diffusion depth to control the extent of influence propagation. Since $\mathbf{S}^{\prime}$ is symmetric, each power $(\mathbf{S}^{\prime})^{k}$ is symmetric, and therefore the diffused matrix $\mathbf{S}^{(d)}$ is also symmetric. The entries $S^{(d)}_{ij}$ represent a score of multi-hop, locality-aware proximity between items $i$ and $j$, learned from the aggregated sequential patterns as illustrated in Figure 3. This diffusion process is analogous to using graph diffusion kernels to define similarity Gasteiger et al. (2019) and converges when $\alpha<1/\rho(\mathbf{S}^{\prime})$. In practice, we estimate $\rho(\mathbf{S}^{\prime})$ (e.g., via power iteration) and choose a suitable $\alpha$.
Lemma 1 (Diffusion Convergence).
The diffusion process defined in Eq. 2 converges as $d\to\infty$ if $\alpha<1/\rho(\mathbf{S}^{\prime})$, where $\rho(\cdot)$ denotes the spectral radius of $\mathbf{S}^{\prime}$.
Figure 3: Multi-hop diffusion encoding item proximity from sequential patterns. Top: $\mathbf{S}^{\prime}$. Middle: Powers $(\mathbf{S}^{\prime})^{k}$. Bottom: $\mathbf{S}^{(3)}$ encoding decayed proximity.
Proof.
The diffusion process as $d\to\infty$ is $\mathbf{S}^{(\infty)}=\sum_{k=1}^{\infty}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$. We can rewrite this as $\mathbf{S}^{(\infty)}=\mathbf{S}^{\prime}\sum_{j=0}^{\infty}(\alpha\mathbf{S}^{\prime})^{j}$ by letting j=k1j=k-1. This series converges if the standard matrix geometric series $\sum_{j=0}^{\infty}M^{j}$, where $M=\alpha\mathbf{S}^{\prime}$, converges. This occurs if the spectral radius $\rho(M)<1$, i.e., $\rho(\alpha\mathbf{S}^{\prime})<1$. Since $\rho(\alpha\mathbf{S}^{\prime})=\alpha\rho(\mathbf{S}^{\prime})$ (for $\alpha>0$), convergence requires $\alpha<1/\rho(\mathbf{S}^{\prime})$. ∎
Symmetric Normalization ($\tilde{\mathbf{S}}$):
The resulting symmetric, diffused matrix $\mathbf{S}^{(d)}$ is then symmetrically normalized to produce the final item-item graph component:
$\tilde{\mathbf{S}}=\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2},\quad\text{where }\mathbf{D}_{S}=\text{diag}(\mathbf{S}^{(d)}\mathbf{1}_{n})$ (3)
This standard normalization technique Chung (1997) yields $\tilde{\mathbf{S}}$, ensuring its weights are scaled appropriately for stable spectral analysis. The weights in $\tilde{\mathbf{S}}$ reflect a transformation of user sequences into measures of item-item sequential relevance.
Interaction Normalization:
The raw user-item interaction matrix $\mathbf{X}$ is normalized to account for varying user activity levels and popularities of the items. Standard normalizations include:
$\tilde{\mathbf{X}}_{U}=\mathbf{D}_{U}^{-1/2}\mathbf{X},\quad\tilde{\mathbf{X}}_{I}=\mathbf{X}\mathbf{D}_{I}^{-1/2}$ (4)
where $\mathbf{D}_{U}=\text{diag}(\mathbf{X}\mathbf{1}_{n})$ and $\mathbf{D}_{I}=\text{diag}(\mathbf{X}^{T}\mathbf{1}_{m})$. For constructing the unified adjacency matrix $\mathbf{A}$, we use the original binary interaction matrix $\mathbf{X}$, as its structure is fundamental. The normalized versions can be used in other components, like the filter design.
Unified Graph Adjacency and Laplacian ($\mathbf{A},\mathbf{L}$):
We integrate the user-item interactions $\mathbf{X}$ and the sequence-derived item-item graph $\tilde{\mathbf{S}}$ into a single, symmetric adjacency matrix for the graph:
$\mathbf{A}=\begin{bmatrix}\mathbf{0}_{m\times m}&\mathbf{X}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}\end{bmatrix}$ (5)
This matrix $\mathbf{A}$ forms a heterogeneous unified graph where the $\mathbf{0}_{m\times m}$ block denotes no direct user-user links are explicitly modeled here, $\mathbf{X}$ captures user-item interactions, and $\tilde{\mathbf{S}}$ encodes the sequence-derived item-item relationships. From $\mathbf{A}$, we compute the symmetric normalized Laplacian, which is the cornerstone of our spectral filtering:
$\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2},\quad\text{where }\mathbf{D}=\text{diag}(\mathbf{A}\mathbf{1})$ (6)
The properties of this Laplacian are discussed in Proposition 2. Our approach, by constructing $\tilde{\mathbf{S}}$ from symmetrized sequential data and then applying diffusion, embeds sequence-derived proximity into a standard spectral framework, offering a distinct way to leverage temporal signals compared to pure sequential models that maintain strict causality Kang & McAuley (2018); Sun et al. (2019).
Proposition 2 (Laplacian Validity).
The Laplacian $\mathbf{L}$, constructed from the symmetric adjacency matrix $\mathbf{A}$, is symmetric, positive semidefinite, and has real eigenvalues in the range $[0,2]$.
Proof sketch; full proof in Appendix B.
This follows from spectral graph theory Chung (1997), as $\mathbf{A}$ is constructed to be symmetric and non-negative, and Eq. 6 defines the standard normalized Laplacian. ∎
3.3 Spectral Filtering Framework
To model user preferences at different levels of specificity, we perform spectral filtering on the Laplacian graph $\mathbf{L}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{T}$ where $\bm{\Lambda}=\text{diag}(\lambda_{1},\ldots,\lambda_{m+n})$ contains eigenvalues in ascending order $0\leq\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{m+n}\leq 2$ and $\mathbf{U}=[\mathbf{u}_{1},\ldots,\mathbf{u}_{m+n}]$ contains the corresponding eigenvectors. In practice, for computational efficiency, we perform a truncated eigendecomposition focusing on $r\ll(m+n)$ which includes the low- and mid-frequency bands relevant to our filters. The lowpass filter $g_{\text{LP}}(\lambda)$ and bandpass filter $g_{\text{BP}}(\lambda)$ operate on the corresponding $r$ eigenpairs by selecting and amplifying the mid-frequency signals present within this subspace.
Proposition 3 (Laplacian Properties).
The eigenvalues of $\mathbf{L}$ lie in $[0,2]$, with the number of zero eigenvalues equal to the number of connected components in the graph.
Bandpass Filtering:
To capture mid-frequency user-specific behavior, we design a bandpass filter using a Gaussian kernel:
$g_{\text{BP}}(\lambda)=\exp\left(-\frac{(\bar{\lambda}-c)^{2}}{w}\right),\quad\bar{\lambda}=\frac{\lambda-\lambda_{\min}}{\lambda_{\max}-\lambda_{\min}}$ (7)
where $c$ and $w$ control the center and width of the frequency band. Our empirical analysis (Appendix G) shows spectral coefficients for intermediate eigenvalues have higher magnitude for user-specific preferences than global trends. This aligns with spectral theory: low frequencies capture global popularity (smooth across graph), high frequencies represent noise (rapid variation), while mid-frequencies are piecewise smooth, which varies between communities but consistent within them. In recommendations, user communities correspond to groups with shared item preferences, making mid-frequency the natural band for personalization.
Let $\mathbf{G}_{\text{BP}}=\text{diag}(g_{\text{BP}}(\lambda_{1}),\ldots,g_{\text{BP}}(\lambda_{r}))$ denote the spectral bandpass filter. We filter the item-degree-normalized user-item interaction signal:
$\mathbf{F}_{BP}=\mathbf{X}\mathbf{D}_{I}^{-1/2}\mathbf{U}\mathbf{G}_{BP}\mathbf{U}^{T}\mathbf{D}_{I}^{-1/2}$ (8)
Lowpass Filtering:
To preserve global structure and smooth preferences, we compute a complementary low-pass component augmented with user-user similarity. To enhance global structure, we augment user-item signals with user-user similarities. We define $\mathbf{X}_{b}=[\mathbf{C}_{U},\mathbf{X}]$, where $\mathbf{C}_{U}=\tilde{\mathbf{X}}_{U}\tilde{\mathbf{X}}_{U}^{T}$. The low-pass component is then computed as:
$\mathbf{F}_{LP}=\mathbf{X}_{b}\mathbf{D}_{b}^{-1/2}\mathbf{U}\mathbf{U}^{T}\mathbf{D}_{b}^{1/2}$ (9)
where $\mathbf{D}_{b}=\text{diag}(\mathbf{X}_{b}^{T}\mathbf{1}_{m})$. We note that $\mathbf{C}_{U}$ is not part of the Laplacian but is introduced to incorporate global user similarity in the lowpass representation. As $\mathbf{X}_{b}$ contains both user-user and user-item components, we extract only the item-related part after filtering by:
$\mathbf{F}_{LP}=\mathbf{F}_{LP}[:,m:]$ (10)
In our framework, mid-frequency components correspond to more personalized patterns, while low frequencies capture smoother, population-level preferences. We compute item scores by fusing mid-frequency personalization with low-frequency global trends. Specifically, we compute:
$\mathbf{Y}=\phi\cdot\mathbf{F}_{\text{BP}}+(1-\phi)\cdot\mathbf{F}_{\text{LP}},\quad\phi\in[0,1]$ (11)
Both $\mathbf{F}_{\text{BP}}$ and $\mathbf{F}_{\text{LP}}$ are outputs of linear spectral filters applied to the graph Laplacian $\mathbf{L}$. Their convex combination remains a valid linear operator, as the space of spectral filters is closed under addition and scalar multiplication Sandryhaila & Moura (2013); Püschel & Moura (2006). Figure 2 illustrates our dual-filter architecture.
[…]
4 Experiments
We evaluate GSPRec to answer: (1) How does sequential graph enrichment affect performance? (2) What is the impact of frequency-aware filtering? (3) How does GSPRec compare to state-of-the-art recommenders?
4.1 Experimental Setup
Data:
We conduct experiments on four benchmark datasets: ML100K, ML1M Harper & Konstan (2015), Netflix Bennett & Lanning (2007), and Amazon Beauty Ni et al. (2019). These datasets maintain comparability with recent spectral methods Xia et al. (2025; 2024). We use an 8:1:1 train/test split with 10% of training data for validation. Statistics are shown in Table 3.
Baselines:
We compare GSPRec with: (i) Popularity; (ii) GCN-based methods: LightGCN He et al. (2020), LR-GCCF Chen et al. (2020b), IMP-GCN Liu et al. (2021), SimpleX Mao et al. (2021a), UltraGCN Mao et al. (2021b); and (iii) GSP-based methods: GF-CF Shen et al. (2021), JGCF Guo et al. (2023), PGSP Liu et al. (2023), HiGSP Xia et al. (2024), FaGSP Xia et al. (2025). Detailed descriptions are in Appendix E. GSPRec leverages temporal graph construction with dual-filter approach targeting both low- and band-pass components.
Implementation Details:
For graph construction, we set the diffusion depth $d=2$ and the decay $\alpha=0.4$ in all datasets to capture immediate and second-order transitions while attenuating longer-range influences. Eigendecomposition dimensionality $r$ varies by dataset size. Filter parameters $c$, $w$, and fusion weight $\phi$ are tuned per dataset as shown in Table 3. Implementation details are in Appendix F.
Table 2: Statistics of evaluation datasets.
Dataset #Users #Items #Inter. Density
ML100K 943 1,682 100,000 6.30%
ML1M 6,040 3,706 1,000,209 4.47%
Netflix 20,000 17,720 5,678,654 1.60%
Beauty 22,363 12,101 198,502 0.07%
Table 3: Filter parameters across datasets.
Dataset $\phi$ $c$ $w$ $r$
ML100K 0.5 0.2 0.1 32
ML1M 0.3 0.8 0.1 128
Netflix 0.3 0.4 0.3 256
Beauty 0.5 0.8 0.3 512
Metrics:
Following prior work He et al. (2020); Shen et al. (2021); Xia et al. (2025); Guo et al. (2023); Xia et al. (2024); Liu et al. (2023), we use Normalized Discounted Cumulative Gain (NDCG@$k$) and Mean Reciprocal Rank (MRR@$k$) at $k\in\{5,10,20\}$, abbreviated as N@k and M@k respectively. We apply spectral filtering to obtain preference scores, exclude previously interacted items, and rank the top-$k$ items.
4.2 Overall Performance
Table 4: Performance comparison on four public datasets. The best performance is denoted in bold. “Improv.” denotes the percentage improvement of GSPRec compared to the strongest baseline method.
Datasets Metric GCN-based Methods GSP-based Methods Improv.
Popularity LightGCN LR-GCCF IMP-GCN SimpleX UltraGCN GF-CF JGCF PGSP FaGSP HiGSP GSPRec
ML100K N@5 0.5672 0.7034 0.5587 0.6764 0.6995 0.6786 0.6875 0.6582 0.6851 0.7106 0.7166 0.7543 +5.26%
M@5 0.5151 0.6362 0.4943 0.6006 0.6397 0.6094 0.6234 0.4471 0.6149 0.6518 0.6629 0.7358 +11.00%
N@10 0.5857 0.6771 0.5603 0.6605 0.6866 0.6688 0.6843 0.6695 0.6722 0.7092 0.7021 0.7572 +6.77%
M@10 0.5296 0.5877 0.4616 0.5690 0.6064 0.5749 0.6010 0.4222 0.5767 0.6384 0.6380 0.7443 +16.59%
N@20 0.5867 0.6618 0.5555 0.6605 0.6543 0.6486 0.6621 0.6730 0.6495 0.6989 0.6724 0.7465 +6.81%
M@20 0.5315 0.5652 0.4411 0.5690 0.5411 0.5289 0.5593 0.4197 0.5496 0.5901 0.5782 0.7457 +26.37%
ML1M N@5 0.0958 0.5845 0.3540 0.5583 0.5934 0.5739 0.5935 0.6121 0.5963 0.6112 0.6062 0.6237 +2.05%
M@5 0.1943 0.5160 0.2997 0.4934 0.5244 0.5087 0.5254 0.4172 0.5313 0.5430 0.5386 0.5715 +5.25%
N@10 0.0960 0.5873 0.3787 0.5594 0.5921 0.5773 0.5897 0.6238 0.5923 0.6082 0.6042 0.6431 +3.10%
M@10 0.2180 0.5010 0.2946 0.4685 0.5051 0.4886 0.4996 0.3869 0.5063 0.5229 0.5148 0.5837 +11.63%
N@20 0.1035 0.5692 0.3980 0.5416 0.5687 0.5600 0.5678 0.6270 0.5687 0.5808 0.5769 0.6431 +2.57%
M@20 0.2309 0.4644 0.2924 0.4292 0.4625 0.4500 0.4557 0.3730 0.4646 0.4748 0.4620 0.5871 +23.65%
Netflix N@5 0.5105 0.6822 0.6023 0.6328 0.6603 0.5680 0.7005 0.7065 0.6811 0.7162 0.7162 0.7201 +0.54%
M@5 0.4505 0.6147 0.5323 0.5528 0.5902 0.4941 0.6360 0.5283 0.6127 0.6524 0.6532 0.6680 +2.27%
N@10 0.5488 0.6814 0.6134 0.6307 0.6694 0.5746 0.6928 0.7148 0.6756 0.7079 0.7062 0.7270 +1.71%
M@10 0.4712 0.5974 0.5234 0.5386 0.5839 0.4708 0.6126 0.4837 0.5905 0.6293 0.6275 0.6762 +7.45%
N@20 0.5604 0.6642 0.6043 0.6152 0.6572 0.5580 0.6689 0.7164 0.6543 0.6814 0.6800 0.7213 +0.68%
M@20 0.4773 0.5636 0.4963 0.4995 0.5555 0.4309 0.5691 0.4273 0.5502 0.5827 0.5800 0.6777 +16.30%
Beauty N@5 0.0149 0.0668 0.0533 0.0570 0.0682 0.0599 0.0659 0.0501 0.0674 0.0712 0.0718 0.0733 +2.09%
M@5 0.0118 0.0530 0.0423 0.0444 0.0544 0.0476 0.0523 0.0535 0.0538 0.0564 0.0574 0.0619 +7.84%
N@10 0.0200 0.0763 0.0635 0.0667 0.0761 0.0649 0.0751 0.0612 0.0765 0.0799 0.0771 0.0879 +10.01%
M@10 0.0139 0.0519 0.0443 0.0451 0.0517 0.0441 0.0517 0.0590 0.0525 0.0549 0.0531 0.0680 +15.25%
N@20 0.0256 0.0829 0.0678 0.0729 0.0819 0.0689 0.0819 0.0721 0.0828 0.0846 0.0799 0.1013 +19.74%
M@20 0.0154 0.0466 0.0388 0.0409 0.0461 0.0393 0.0469 0.0622 0.0481 0.0485 0.0456 0.0718 +15.43%
Table 4 presents performance comparisons. GSPRec consistently outperforms all baselines across metrics and datasets, with improvements over recent GSP-based methods validating the effectiveness of leveraging sequential graph construction with mid-frequency spectral filtering.
Performance improvements range from 0.54% to 26.37% across datasets and metrics. Beauty dataset achieves 19.74% improvement on N@20, while ML100K shows improvements ranging from 5.26% to 26.37% and Netflix demonstrates gains of 0.54% to 16.30%. This variation indicates that method effectiveness depends on dataset-specific characteristics that extend beyond simple density metrics. The performance variation correlates with dataset-specific optimal parameter configurations shown at Table 3. ML100K exhibits optimal performance at lower frequency centers, Netflix at mid-range frequencies, while Beauty and ML1M favor higher frequencies. These indicate that GSPRec adapts to the underlying spectral structure of each dataset rather than imposing a fixed frequency range.
4.3 Ablation Analysis
Table 5: Ablation study on ML1M dataset.
Variant N@10 $\pm$ SE M@10 $\pm$ SE
GSPRec-NB 0.5769 0.0082 0.5148 0.0093
GSPRec-NL 0.6042 0.0079 0.5229 0.0088
GSPRec-NS 0.6274 0.0039 0.5704 0.0050
GSPRec-SE 0.6416 0.0037 0.5831 0.0049
GSPRec 0.6431 0.0074 0.5837 0.0081
Table 6: Filtering comparison on ML1M dataset.
Components N@10 $\pm$ SE M@10 $\pm$ SE
$\mathbf{F}_{\text{LP}}$ only 0.5769 0.0082 0.5148 0.0093
$\mathbf{F}_{\text{BP}}$ only 0.6042 0.0079 0.5229 0.0088
$\mathbf{F}_{\text{LP}}$$\mathbf{F}_{\text{BP}}$ ($\phi$=0.3) 0.6431 0.0074 0.5837 0.0081
$\mathbf{F}_{\text{LP}}$$\mathbf{F}_{\text{BP}}$ ($\phi$=0.5) 0.6327 0.0076 0.5714 0.0085
$\mathbf{F}_{\text{LP}}$$\mathbf{F}_{\text{BP}}$ ($\phi$=0.7) 0.6281 0.0078 0.5682 0.0087
Table 6 reveals a clear component hierarchy. Removing bandpass filtering causes the largest performance degradation, followed by lowpass filtering removal, while sequential diffusion removal causes the smallest drop. This hierarchy demonstrates that spectral filtering strategy contributes more substantially to performance than graph augmentation through sequential information. Table 6 confirms that bandpass filtering alone outperforms lowpass alone by 4.7%, supporting the effectiveness of mid-frequency targeting over traditional low-frequency approaches.
Parameter analysis reveals both universal architectural constants and dataset-specific adaptations. Diffusion parameters $\alpha=0.4$ and $d=2$ remain across all datasets, which suggests fundamental properties of sequential transition structures that transcend domain-specific characteristics. However, spectral parameters exhibit systematic variation that eigenspace dimensionality $r$ scales linearly with dataset size, while bandpass center positions cluster according to dataset characteristics, in which indicating that different datasets exhibit distinct spectral signatures in their interaction patterns.
The fusion weight $\phi$ provides interpretable insight into the balance between personalization and popularity signals. The optimal value of 0.3 on ML1M indicates that bandpass-based personalization dominates the recommendation process while lowpass popularity signals provide complementary information. This quantifies the relative importance of different spectral components and demonstrates how GSPRec achieves effective signal separation.
4.4 Parameter Sensitivity
Figure 4: Effect of bandpass center position ($c$) on NDCG@10 across datasets. Each line represents a different filter bandwidth ($w$).
Figure 5: Effect of bandpass width ($w$) on NDCG@10 across datasets. Each line represents a different center position ($c$). Optimal width depends on center position and dataset characteristics.
Figures 4 and 5 demonstrate that optimal filter configurations depend on both center position and dataset characteristics. The interaction between center and width parameters confirms that different datasets require distinct spectral windows to capture personalized patterns while suppressing noise amplification. This adaptive behavior explains why fixed low-pass or high-pass approaches fail to achieve consistent improvements across diverse recommendation scenarios.
Notably, the sensitivity curves exhibit relative flatness around optimal parameter values, indicating that moderate parameter deviations do not substantially degrade performance. This robustness suggests that GSPRec does not require precise hyperparameter tuning.
4.5 Computational Analysis
Table 7 shows that GSPRec requires 1.27 minutes on ML1M, comparable to other GSP methods (1.18-2.36 minutes) and substantially more efficient than GCN approaches (16-135 minutes). This positions GSPRec on the accuracy-efficiency frontier, unlike GCN methods that achieve moderate improvements at high computational cost, or GSP methods that sacrifice accuracy for speed, GSPRec provides both computational efficiency and high accuracy by targeting the most informative bands.
Table 7: Runtime comparison on ML1M dataset (minutes).
Method LightGCN SimpleX UltraGCN GF-CF PGSP FaGSP HiGSP GSPRec
Runtime 135.96 109.83 16.09 1.18 2.36 2.08 2.04 1.27
5 Conclusion
We presented GSPRec, a temporal-spectral collaborative filtering framework that integrates sequential transitions with dual-frequency spectral filtering. our approach consistently outperforms baselines across four datasets by targeting mid-frequency components, with bandpass filtering achieving 4.7% improvement over state-of-the-art methods.
Limitations and Future Directions: The symmetric construction trades directional information for spectral stability, and eigendecomposition creates computational bottlenecks addressable through polynomial filtering. This work demonstrates that restructuring graph spectra through sequential encoding enables more effective personalization than filtering static user-item graphs.
Appendices
Appendix A Table of Notations
We summarize the key notations used throughout this paper in Table 8.
Table 8: Summary of Notations
Notation Description
Input Data Representations
$\mathcal{U},\mathcal{I}$ Set of users and items, respectively
$\mathcal{D}$ Historical interactions as $(u,i,t)$ triplets
$\mathbf{X}\in\{0,1\}^{m\times n}$ Binary user-item interaction matrix
$\mathcal{S}_{u}$ Time-ordered sequence of items for user $u$
Graph Construction
$\mathbf{S}$ Initial directed item transition matrix from $\mathcal{S}_{u}$
$\mathbf{S}^{\prime}$ Symmetric item connectivity graph derived from $\mathbf{S}$
$\alpha,d$ Diffusion decay factor and depth
$\mathbf{S}^{(d)}$ Multi-hop diffusion matrix from $\mathbf{S}^{\prime}$ up to depth $d$
$\tilde{\mathbf{S}}$ Normalized diffusion matrix derived from $\mathbf{S}^{(d)}$
$\mathbf{A}$ Unified user-item adjacency matrix
$\mathbf{L}$ Normalized Laplacian matrix
Spectral Filtering
$\mathbf{U},\bm{\Lambda}$ Matrices of eigenvectors and eigenvalues of $\mathbf{L}$, respectively
$r$ Number of eigenvectors in truncated decomposition
$g_{\text{BP}}(\lambda)$ Gaussian bandpass filter kernel
$c,w$ Center and width of bandpass filter
$\mathbf{F}_{\text{BP}}$ Bandpass filtered user-item signal
$\mathbf{F}_{\text{LP}}$ Low-pass filtered user-item signal
$\phi$ Fusion parameter balancing filter components
$\mathbf{Y}$ Final prediction matrix
Appendix B Algorithmic Details
B.1 GSPRec Algorithm
The full end-to-end procedure of our method is outlined in Algorithm 1. It summarizes the construction of the sequential graph encoding, Laplacian computation, spectral filtering, and final score generation.
Input: Interaction set $\mathcal{D}=\{(u,i,t)\}$, diffusion order $d$, decay factor $\alpha$, filter parameters $(c,w)$, fusion weight $\phi$
Output: Predicted score matrix $\mathbf{Y}\in\mathbb{R}^{m\times n}$
1ex// Graph Construction
$\mathbf{X}\leftarrow$ Binary user-item matrix from $\mathcal{D}$
$\mathcal{S}_{u}\leftarrow$ Chronologically sorted item sequences per user
$\mathbf{S}\leftarrow$ Initial directed item transition matrix from $\mathcal{S}_{u}$ (where $S_{ij}=1$ if $i\to j$)
[…]
$\mathbf{S}^{(d)}\leftarrow\sum_{k=1}^{d}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$ // Exponential diffusion on symmetric $\mathbf{S}^{\prime}$
$\tilde{\mathbf{S}}\leftarrow\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2}$ // where $\mathbf{D}_{S}=\text{diag}(\mathbf{S}^{(d)}\mathbf{1}_{n})$
$\mathbf{A}\leftarrow\begin{bmatrix}\mathbf{0}&\mathbf{X}\\ \mathbf{X}^{\top}&\tilde{\mathbf{S}}\end{bmatrix}$
$\mathbf{D}\leftarrow\text{diag}(\mathbf{A}\mathbf{1})$
$\mathbf{L}\leftarrow\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$
$\mathbf{U},\mathbf{\Lambda}]\leftarrow$ Eigendecomposition of $\mathbf{L}$
// Frequency-Aware Spectral Filtering
$\bar{\bm{\lambda}}\leftarrow$ Diagonal entries of $\mathbf{\Lambda}$
$\mathbf{G}_{\text{BP}}\leftarrow\text{diag}\left(\exp\left(-\frac{(\bar{\bm{\lambda}}-c)^{2}}{w}\right)\right)$ // Gaussian band-pass
$\mathbf{F}_{\text{BP}}\leftarrow\mathbf{X}\mathbf{D}_{I}^{-1/2}\mathbf{U}\mathbf{G}_{\text{BP}}\mathbf{U}^{\top}\mathbf{D}_{I}^{-1/2}$
// Low-Pass Filtering
$\tilde{\mathbf{X}}_{U}\leftarrow\mathbf{D}_{U}^{-1/2}\mathbf{X}$ // where $\mathbf{D}_{U}=\text{diag}(\mathbf{X1}_{n})$, $\mathbf{D}_{I}=\text{diag}(\mathbf{X}^{T}\mathbf{1}_{m})$
$\mathbf{C}_{U}\leftarrow\tilde{\mathbf{X}}_{U}\tilde{\mathbf{X}}_{U}^{\top}$
$\mathbf{X}_{b}\leftarrow[\mathbf{C}_{U},\mathbf{X}]$
$\mathbf{D}_{b}\leftarrow\text{diag}(\mathbf{X}_{b}^{\top}\mathbf{1})$
$\mathbf{F}_{\text{LP}}\leftarrow\mathbf{X}_{b}\mathbf{D}_{b}^{-1/2}\mathbf{U}\mathbf{U}^{\top}\mathbf{D}_{b}^{1/2}$
$\mathbf{F}_{\text{LP}}\leftarrow\mathbf{F}_{\text{LP}}[:,m{:}]$ // Select only item columns
// Fusion and Output
$\mathbf{Y}\leftarrow\phi\cdot\mathbf{F}_{\text{BP}}+(1-\phi)\cdot\mathbf{F}_{\text{LP}}$
return $\mathbf{Y}$
Algorithm 1 GSPRec
B.2 Proof of Proposition 2: Laplacian Validity
Proof.
We first establish properties of the adjacency matrix $\mathbf{A}\in\mathbb{R}^{(m+n)\times(m+n)}$, defined as:
$\mathbf{A}=\begin{pmatrix}\mathbf{0}_{m\times m}&\mathbf{X}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}\end{pmatrix}$ (12)
where $\mathbf{X}\in\{0,1\}^{m\times n}$ (the binary user-item interaction matrix) and $\tilde{\mathbf{S}}=\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2}\in\mathbb{R}^{n\times n}$ (the normalized sequence-derived item-item graph, with $\mathbf{S}^{(d)}$ being the diffused matrix from Eq. 2).
Lemma 4.
The adjacency matrix $\mathbf{A}$ is symmetric and non-negative.
Proof of Lemma 4.
To prove symmetry of $\mathbf{A}$, we note:
$\mathbf{A}^{T}$ $=\begin{pmatrix}\mathbf{0}_{m\times m}^{T}&(\mathbf{X}^{T})^{T}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}^{T}\end{pmatrix}$ (13)
$=\begin{pmatrix}\mathbf{0}_{m\times m}&\mathbf{X}\\ \mathbf{X}^{T}&\tilde{\mathbf{S}}\end{pmatrix}=\mathbf{A}$ (14)
This equality holds because $\mathbf{0}_{m\times m}$ is symmetric, $(\mathbf{X}^{T})^{T}=\mathbf{X}$, and $\tilde{\mathbf{S}}^{T}=\tilde{\mathbf{S}}$. The matrix $\tilde{\mathbf{S}}$ is symmetric by construction: the base matrix $\mathbf{S}^{\prime}$ used in the diffusion is symmetric, thus $\mathbf{S}^{(d)}$ (from Eq. 2) is symmetric, and the symmetric normalization (Eq. 3) preserves this symmetry.
For non-negativity of $\mathbf{A}$, we show all its elements are non-negative:
  • $\mathbf{X}\in\{0,1\}^{m\times n}$ by definition, so its elements are $\geq 0$. Thus, $\mathbf{X}\geq 0$ and $\mathbf{X}^{T}\geq 0$.
  • For $\tilde{\mathbf{S}}$: The initial directed transition matrix $\mathbf{S}$ has $s_{ij}\in\{0,1\}$, so $\mathbf{S}\geq 0$. The symmetric base graph $\mathbf{S}^{\prime}$ is constructed from $\mathbf{S}$ such that its entries $s^{\prime}_{ij}\in\{0,1\}$, so $\mathbf{S}^{\prime}\geq 0$. Given $\mathbf{S}^{\prime}\geq 0$ and the decay factor $\alpha\in(0,1)>0$, each power $(\mathbf{S}^{\prime})^{k}$ in the diffusion sum will have non-negative entries. Therefore, $\mathbf{S}^{(d)}=\sum_{k=1}^{d}\alpha^{k-1}(\mathbf{S}^{\prime})^{k}$ (Eq. 2) will also have non-negative entries, i.e., $\mathbf{S}^{(d)}\geq 0$. The degree matrix $\mathbf{D}_{S}=\text{diag}(\mathbf{S}^{(d)}\mathbf{1}_{n})$ will have non-negative diagonal entries. For non-isolated nodes in the graph represented by $\mathbf{S}^{(d)}$, these entries will be positive, making $\mathbf{D}_{S}^{-1/2}$ well-defined with non-negative real entries. Hence, $\tilde{\mathbf{S}}=\mathbf{D}_{S}^{-1/2}\mathbf{S}^{(d)}\mathbf{D}_{S}^{-1/2}\geq 0$.
Therefore, $\mathbf{A}$ is both symmetric and non-negative. ∎
Given that $\mathbf{A}$ is symmetric and non-negative (by Lemma 4), and its degree matrix $\mathbf{D}=\text{diag}(\mathbf{A}\mathbf{1})$, the normalized Laplacian is $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ (Eq. 6). We apply Theorem 5:
Theorem 5 (Properties of Normalized Laplacian (Chung, 1997 Chung (1997))).
For a symmetric, non-negative adjacency matrix $\mathbf{A}$ with degree matrix $\mathbf{D}$, the normalized Laplacian $\mathbf{L}=\mathbf{I}-\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ satisfies: (1) $\mathbf{L}$ is real and symmetric. (2) $\mathbf{L}$ is positive semi-definite. (3) All eigenvalues of $\mathbf{L}$ lie in $[0,2]$.
Applying Theorem 5 directly to our $\mathbf{A}$ and $\mathbf{L}$ concludes that $\mathbf{L}$ is symmetric, positive semi-definite, and has eigenvalues in $[0,2]$. ∎
Appendix C Theoretical Analysis of Spectral Bands
The concentration of user-specific patterns in mid-frequency components can be understood through the lens of graph signal smoothness. Low-frequency components correspond to signals that are smooth across the graph, capturing global popularity trends that vary slowly across the user-item structure. High-frequency components often represent noise or sparse outlier preferences.
Mid-frequency components represent a balance where signals vary moderately across local graph neighborhoods - precisely the pattern exhibited by personalized preferences that differ between user communities but show consistency within them. This aligns with previous findings in spectral clustering Von Luxburg (2007) and network community detection Newman (2006), where mid-frequency eigenvectors have been shown to encode community structure.
We can formalize this intuition through the concept of total variation of a signal Shuman et al. (2013). Given a signal $\mathbf{f}$ on a graph with Laplacian $\mathbf{L}$, the quadratic form $\mathbf{f}^{T}\mathbf{L}\mathbf{f}$ measures how smoothly the signal varies across the graph. For the eigenvectors $\mathbf{u}_{i}$ of $\mathbf{L}$, we have $\mathbf{u}_{i}^{T}\mathbf{L}\mathbf{u}_{i}=\lambda_{i}$, where $\lambda_{i}$ is the corresponding eigenvalue.
User preference signals can be decomposed into these eigenvectors:
$\mathbf{f}=\sum_{i=1}^{m+n}\alpha_{i}\mathbf{u}_{i}$ (15)
In our empirical analysis across datasets, we consistently observe that coefficients $\alpha_{i}$ corresponding to mid-frequency components ($0.3<\lambda_{i}<0.8$) have significantly higher magnitude for user-specific preference patterns compared to global trends. This provides empirical validation of our theoretical understanding of spectral representation in recommendation graphs.
Appendix D Computational Complexity Analysis
GSPRec has time complexity dominated by eigendecomposition and filtering operations. For sparse datasets with $r\ll(m+n)$, the method remains tractable on benchmark datasets as demonstrated by our runtime analysis (Table 7). For large-scale deployments, approximation techniques like polynomial filtering can reduce computational requirements to $O(K\cdot|\mathcal{D}|)$ for $K$-order polynomials, making the approach practical for industrial applications.
Appendix E Baseline Details
We compare GSPRec with a comprehensive set of baselines representing both GCN-based and GSP-based CF methods:
GCN-based CF Methods:
  • LightGCN He et al. (2020) simplifies GCN for recommendation by removing feature transformation and non-linear activation, retaining only essential neighborhood aggregation.
  • LR-GCCF Chen et al. (2020b) introduces residual connections to graph convolutional networks for improving recommendation performance.
  • IMP-GCN Liu et al. (2021) propagates information in user sub-graphs to reduce the impact of noise and alleviate the over-smoothing problem.
  • SimpleX Mao et al. (2021a) focuses on optimizing loss function and negative sampling strategies.
  • UltraGCN Mao et al. (2021b) approximates infinite-layer graph convolutions via constraint loss and flexible edge weight assignment.
GSP-based CF Methods:
  • GF-CF Shen et al. (2021) integrates linear and ideal low-pass filters to model user preferences.
  • JGCF Guo et al. (2023) uses Jacobi polynomial-based filters to amplify high-frequency signals that are typically suppressed by standard GCNs.
  • PGSP Liu et al. (2023) uses a mixed-frequency low-pass filter over personalized graph signals.
  • HiGSP Xia et al. (2024) adopts cluster-wise and globally-aware filters to recognize unique and general interaction patterns.
  • FaGSP Xia et al. (2025) employs Cascaded and Parallel Filter Modules for user preference modeling.
For comprehensive evaluation, We use official repositories for methods for reproducing results on our dataset settings. We report results for LR-GCCF, IMP-GCN, SimpleX, UltraGCN, and GF-CF as presented in Xia et al. (2024; 2025), which used identical dataset preprocessing and evaluation protocols.
E.1 Filter Design
Figure 6: Frequency response curves of different filter configurations all datasets. The low-pass filter (blue) preserves signals in the lower 30% of the spectrum. The band-pass filter (green) selectively amplifies mid-frequency components associated with personalized patterns. The combined filter response (purple) illustrates how dual-filter captures complementary signals across the spectrum.
Appendix F Reproducibility
To improve reproducibility, we provide the complete algorithm of GSPRec in Algorithm 1, which details the graph construction, spectral decomposition, and dual filtering processes. The optimal hyperparameters for all experimental datasets are summarized in Table 3. For implementation, we process datasets following the procedure described in Section 4.1, which involves splitting datasets in an 8:2 ratio for training/testing, with 10% of the training set used for validation. When implementing GSPRec, the key parameter settings are:
  • Diffusion depth $d=2$ and decay factor $\alpha=0.4$ for sequential graph construction
  • Dataset-specific eigenvector count $r$, band-pass center $c$, width $w$, and fusion weight $\phi$ as specified in Table 3
For evaluation, we follow the protocol detailed in Section 4.1, using NDCG and MRR metrics at $k\in\{5,10,20\}$ after excluding previously interacted items.
F.1 Computational Resources
Experiments were conducted on a system with AMD EPYC processors and NVIDIA RTX A6000 GPUs. The computational requirements varied with dataset size and the number of eigenvectors used. The eigendecomposition step is the primary computational bottleneck as identified in our complexity analysis (Appendix D). For benchmark datasets used in our experiments, the runtime and memory requirements were reasonable, allowing for comprehensive evaluation on standard research hardware. Our experimental protocol, including parameter tuning and ablation studies, involved multiple runs for each configuration to ensure statistical significance.
F.2 Scaling to Large-Scale Systems
For industrial-scale applications with millions of users and items, we propose three optimization strategies:
F.2.1 Approximate Eigendecomposition
The dominant computational bottleneck can be addressed using the Nyström method Fowlkes et al. (2004) or randomized SVD approaches Halko et al. (2011), reducing complexity from $O(r(m+n)^{2})$ to $O(r^{2}(m+n))$ (or better, depending on specific variant, e.g., linear in $(m+n)$ for matrix-vector products), with minimal performance impact.
F.2.2 Graph Sparsification
For large-scale graphs, we can apply spectral sparsification techniques Spielman & Teng (2011) that preserve the essential spectral properties while reducing edge density, accelerating both graph construction and eigendecomposition.
F.2.3 Incremental Updates
In production environments, the full spectral decomposition need not be recomputed for each recommendation. We propose an incremental update scheme where the spectral components are updated periodically while incremental preference updates can be applied in real-time.
Appendix G Mathematical Analysis of Bandpass Filtering
We provide mathematical justification for our bandpass filtering approach and its relevance to recommendation systems. Our Gaussian kernel on the graph Laplacian eigenspectrum is defined as in Equation 7.
$g(\lambda)=\exp\left(-\frac{(\bar{\lambda}-c)^{2}}{w}\right),\quad\bar{\lambda}=\frac{\lambda-\lambda_{\min}}{\lambda_{\max}-\lambda_{\min}}$
This satisfies the formal definition of a bandpass filter: it achieves maximum response when $\bar{\lambda}=c$ (which corresponds to $\lambda_{c}=c(\lambda_{\max}-\lambda_{\min})+\lambda_{\min}$ in the original eigenvalue domain) and attenuates components with eigenvalues far from $\lambda_{c}$. For our normalized Laplacian with original eigenvalues $\lambda\in[0,2]$ (so $\lambda_{\min}\approx 0,\lambda_{\max}\approx 2$), setting $c\in(0,1)$ (the center in the normalized $\bar{\lambda}\in[0,1]$ domain) ensures the filter passes a band of mid-range eigenvalues while attenuating both low and high extremes. When applied to user preference signals decomposed as $\mathbf{x}=\sum_{i=1}^{m+n}\alpha_{i}\mathbf{u}_{i}$, the filtered output $\mathbf{y}=\mathbf{U}\text{diag}(g(\bm{\Lambda}))\mathbf{U}^{T}\mathbf{x}$ selectively preserves components $\alpha_{i}\mathbf{u}_{i}$ where the eigenvalue $\lambda_{i}$ (corresponding to $\mathbf{u}_{i}$) results in a high $g(\lambda_{i})$ value.
Our empirical analysis across recommendation datasets reveals that coefficients $\alpha_{i}$ corresponding to intermediate eigenvalues ($0.3<\lambda_{i}<0.8$, assuming $\lambda_{i}$ are original eigenvalues in $[0,2]$) have higher magnitude for user-specific preference patterns compared to global trends. This confirms that mid-frequency components encode valuable personalization signals that traditional low-pass approaches would suppress.
Table 6 provides quantitative evidence of this phenomenon: while low-pass filtering captures popularity trends, our bandpass approach more effectively preserves the user-specific patterns essential for accurate personalized recommendations.